### v1

# reference.bib

Abstract.
[…]
Acknowledgements.
The author was supported by the HIMR/UKRI "Additional Funding Programme for Mathematical Sciences", grant number EP/V521917/1 as well as by a scholarship by Onassis Foundation. The author would also like to thank Prof. Thanasis Bouganis for suggesting this problem as well as for his continuous guidance and support.
1. Introduction
Kohnen and Skoruppa in [kohnen_skoruppa] considered the Dirichlet series
$\mathcal{D}_{F,G}(s)=\sum_{N=1}^{\infty}\langle\phi_{N},\psi_{N}\rangle N^{-s},$
[…]
$\mathcal{D}_{F,G}(s)=\langle\phi_{1},\psi_{1}\rangle Z_{F}(s),$
where $Z_{F}(s)$ is the spinor $L-$function attached to $F$. This gives a way of proving the meromorphic continuation of $Z_{F}(s)$ to $\mathbb{C}$.
[…]
In this paper, we consider a Dirichlet series of the same type, for two cups forms $F,G$ of orthogonal groups of real signature $(2,n+2)$, which are eigenforms for their corresponding Hecke algebras. Here, we take $G$ to be a Maass lift of a Poincaré series. The Maass space in this setting has been defined by Gritsenko in [gritsenko_jacobi], Sugano in [sugano_lifting] and Krieg in [krieg_jacobi]. Our main goal is to show how the method of Kohnen and Skoruppa in [kohnen_skoruppa] can be extended in this setting and investigate the connection of this Dirichlet series with the standard $L-$function attached to $F$. The question of its analytic continuation and functional equation is quite different in nature and will be the subject matter of an upcoming paper.
There is a variety of reasons why we should expect such a connection. First of all, the orthogonal groups of signature $(2,n+2)$ are isomorphic to classical groups for small $n$. For example, if $n=1$, there is an isomorphism with the symplectic and paramodular groups of degree 22 22. For $n=2$, there is an isomorphism with the Hermitian groups of degree 22 22, for $n=4$ with quaternionic groups of degree 22 22 and for $n=6$ with the linear groups of Cayley numbers of degree 22 22 (see for example [eisenstein_thesis, page 5]). Therefore, already the results of Kohnen and Skoruppa in [kohnen_skoruppa] and Gritsenko in [gritsenko] should indicate a connection in the $n=1,2$ case. Moreover, it should be noted that Gritsenko in [gritsenko_dirichlet_orthogonal] initiated the study of a Dirichlet series involving the Fourier coefficients of an orthogonal modular form and its connection with the standard $L-$function attached to it, following factorisation methods similar to [gritsenko]. Sugano in [sugano] extended Gritsenko’s result, proving an Euler product relation for a Dirichlet series involving twists of the Fourier coefficients by functions on an orthogonal group of lower rank. His work can be seen as an extension of Andrianov’s work in [andrianov] in the orthogonal setting. As we mentioned, our methods in this paper can be considered as an extension of the methods employed by Kohnen and Skoruppa in [kohnen_skoruppa] and therefore the result in [sugano] by Sugano is pivotal. Finally, results proved by Shimura in [shimura2004arithmetic] on the sets of solutions $\phi[x]=q$, where $\phi$ is a bilinear form and $q\in\mathbb{Q}^{\times}$, turn out to be crucial.
Let us now give a brief outline of our results. The setup is as follows: We start with an even positive definite symmetric matrix $S$ of rank $n\geq 1$ so that the lattice $L:=\mathbb{Z}^{n}$ is maximal. We then set
$S_{0}=\begin{pmatrix}&&1\\ &-S&\\ 1&&\end{pmatrix},S_{1}=\begin{pmatrix}&&1\\ &S_{0}&\\ 1&&\end{pmatrix}$
of real signature $(1,n+1)$ and $(2,n+2)$ respectively. If now $K$ is a field containing $\mathbb{Q}$, we define the corresponding special orthogonal groups of $K-$rational points via
$G^{*}_{K}=\{g\in SL_{n+2}(K)\mid g^{t}S_{0}g=S_{0}\},G_{K}=\{g\in SL_{n+4}(K)\mid g^{t}S_{1}g=S_{1}\}.$
It turns out that $G_{\mathbb{R}}^{0}$, the connected component of the identity, acts on a suitable upper half plane, which we will call $\mathcal{H}_{S}\subset\mathbb{C}^{n+2}$, and a suitable factor of automorphy $j(\gamma,Z)$ can be defined. Let us now denote by $\Gamma_{S}$ the intersection of $G_{\mathbb{R}}^{0}$ with the stabiliser of the lattice $L_{1}=\mathbb{Z}^{n+4}$ and let $F:\mathcal{H}_{S}\longrightarrow\mathbb{C}$ be an orthogonal cusp form with respect to $\Gamma_{S}$ (see Definitions 2.2 and 2.5).
If now $Z=(\tau^{\prime},z,\tau)\in\mathcal{H}_{S}$, we can write
$F(Z)=\sum_{m\geq 1}\phi_{m}(\tau,z)e^{2\pi im\tau^{\prime}},$
[…]
$\mathcal{D}_{F,G}(s)=\sum_{m\geq 1}\langle\phi_{m},\psi_{m}\rangle m^{-s},$
[…]
$\mathcal{P}_{k,D,r}(\tau^{\prime},z,\tau):=\sum_{N\geq 1}V_{N}P_{k,D,r}(\tau,z)e(N\tau^{\prime}),$
[…]
$H(\xi)_{\mathbb{Q}}=\{g\in G_{\mathbb{Q}}^{*}\mid g\xi=\xi\}.$
Consider then the case when the class number $h:=\#(H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$ equals 11 11, where $C=\{x\in G_{\mathbb{A}}^{*}\mid xL_{0}=L_{0}\}$. Here, the subscript A denotes the adelization of an algebraic linear group (see Section 6). Then, our main theorem 7.2 reads:
Theorem 1.1.
With the notation and assumptions as above, let $\mathcal{P}$ denote a sufficiently large finite set of prime numbers. Define then the Dirichlet series
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}\langle\phi_{N},V_{N}P_{k,D,r}\rangle N^{-s}.$
We then have that $L_{\mathcal{P}}\left(\mathbf{1};s-k+\frac{n+3}{2}\right)\mathcal{D}_{F,P_{k,D,r},\mathcal{P}}(s)$ can be written as:
$A(\xi)L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\zeta_{\xi,\mathcal{P}}(s-k+n+1)\times\begin{cases}1&\textup{if }n\textup{ odd}\\ \zeta_{\mathcal{P}}(2s-2k+n+2)^{-1}&\textup{if }n\textup{ even}\end{cases}.$
Here, $L(-,s)$ denotes the standard $L-$function attached to orthogonal modular forms and $\mathbf{1}$ denotes the constant function on $H(\xi)_{\mathbb{A}}$. Also $\zeta_{\xi}(s)$ is a Dirichlet series counting the number of solutions of congruences (see Proposition 6.3). For any zeta function (or Dirichlet series) appearing, we use the subscript $\mathcal{P}$ to denote that we don’t take into account the $p-$factors (or $p-$terms) for $p\in\mathcal{P}$.
Note here that the Theorem 7.2 covers the case $h>1$ as well. The rest of the paper (Section 8) is devoted in obtaining explicit examples where $h=1$. In those cases, we obtain a clear-cut Euler product connection between the specific Dirichlet series $\mathcal{D}_{F,\mathcal{P}_{k,-q,r},\mathcal{P}}$ and $L_{\mathcal{P}}(F,s)$ (here we make a particular choice for $D$, i.e. $D=-q$).
We find all instances when this happens in the rank 11 11 case and we give the result in Theorem 8.1. Note that Theorem 8.1 recovers the case of Kohnen and Skoruppa in [kohnen_skoruppa] (we take $S=2$). What is more, it already produces new results related to $L-$functions attached to paramodular forms, as it is not clear how can one use the methods in [kohnen_skoruppa] to produce such results for paramodular forms directly. Finally, we provide at least one explicit example for each of the cases $n=2,4,6,8$, giving us Theorem 8.2. These examples can be related to other classical modular forms, such as Hermitian ($n=2$) and Quaternionic ($n=4$).
[…]
Notation.
Let $\mathbb{Z},\mathbb{Q},\mathbb{R},\mathbb{C}$ denote the usual rings of integers, rationals, real and complex numbers, respectively. For a complex number $z$, we denote by $e(z):=e^{2\pi iz}$. We will use the bracket notation $A[B]:=\overline{B}^{t}AB$ for suitably sized complex matrices $A,B$.
2. Preliminaries
Let $V$ denote a finite dimensional vector space over $\mathbb{Q}$. We start with the following definition:
Definition 2.1.
A $\mathbb{Z}-$lattice $\Lambda$ is a free, finitely generated $\mathbb{Z}-$module, which spans $V$ over $\mathbb{Q}$, together with a symmetric non-degenerate bilinear form $V\times V\longrightarrow\mathbb{Q}$.
In the following, let $V:=\mathbb{Q}^{n}$ and $L:=\mathbb{Z}^{n}$ with $n\geq 1$. Assume $S$ is an even integral positive definite symmetric matrix of rank $n\geq 1$. Even here means $S[x]\in 2\mathbb{Z}$ for all $x\in L$.
Let then $V_{0}:=\mathbb{Q}^{n+2}$ and consider
$S_{0}=\begin{pmatrix}0&0&1\\ 0&-S&0\\ 1&0&0\end{pmatrix}$
of real signature $(1,n+1)$. Define also the bilinear form
$\displaystyle\phi_{0}:V_{0}\times V_{0}$ $\displaystyle\longmapsto\mathbb{Q}$
$\displaystyle(x,y)$ $\displaystyle\longmapsto\frac{1}{2}x^{t}S_{0}y.$
Consider now a Witt decomposition of the form
$V_{0}=e_{n+1}\perp V\perp e_{1},$
where $e_{i}$ is the $i$’th standard basis vector. We then have that $\phi:=\phi\mid_{V\times V}$ is just $(x,y)\longmapsto-\frac{1}{2}x^{t}Sy$, and we make the assumption that $L=\mathbb{Z}^{n}$ is a maximal $\mathbb{Z}-$lattice with respect to $\phi$. Here, maximal means that the lattice $L$ is maximal among all $\mathbb{Z}-$lattices in $V$ on which $\phi[x]$ takes values in $\mathbb{Z}$, where we use the notation $\phi[x]:=\phi(x,x)$ for $x\in V$.
From [shimura2004arithmetic, Lemma 6.3], we then obtain that $L_{0}:=\mathbb{Z}^{n+2}$ is a $\mathbb{Z}-$maximal lattice in $V_{0}=\mathbb{Q}^{n+2}$.
Moreover, define $S_{1}=\begin{pmatrix}&&1\\ &S_{0}&\\ 1&&\end{pmatrix}$ and define the corresponding special orthogonal groups:
$G^{*}_{\mathbb{Q}}:=\{g\in\textup{SL}_{n+2}(\mathbb{Q})\mid g^{t}S_{0}g=S_{0}\},$
$G_{\mathbb{Q}}:=\{g\in\textup{SL}_{n+4}(\mathbb{Q})\mid g^{t}S_{1}g=S_{1}\}.$
We view $G^{*}_{\mathbb{Q}}$ as a subgroup of $G_{\mathbb{Q}}$ via the embedding $g\longmapsto\begin{pmatrix}1&&\\ &g&\\ &&1\end{pmatrix}$.
We also use the notation $SO^{\phi_{0}}(V_{0})$ and in general $SO^{\phi}(V)$ for any quadratic space $(V,\phi)$.
Let now $\mathcal{H}_{S}$ denote one of the connected components of
$\{Z\in V_{0}\otimes_{\mathbb{Q}}\mathbb{C}\mid\phi_{0}[\textup{Im}Z]>0\}.$
In particular, if we denote by $\mathcal{P}_{S}:=\{y=(y_{1},y,y_{2})\in\mathbb{R}^{n+2}\mid y_{1}>0,\phi_{0}[y]>0\}$, we choose:
$\mathcal{H}_{S}=\{z=x+iy\in V_{0}\otimes_{\mathbb{R}}\mathbb{C}\mid y\in\mathcal{P}_{S}\}.$
For a matrix $g\in\textup{Mat}(n+4,\mathbb{R})$, we write it as
$g=\begin{pmatrix}\alpha&a^{t}&\beta\\ b&A&c\\ \gamma&d^{t}&\delta\end{pmatrix},$
with $A\in\textup{Mat}(n+2,\mathbb{R}),\alpha,\beta,\gamma,\delta\in\mathbb{R}$ and $a,b,c,d$ real column vectors. Now the map
$Z\longmapsto g\langle Z\rangle=\frac{-\frac{1}{2}S_{0}[Z]b+AZ+c}{-\frac{1}{2}S_{0}[Z]\gamma+d^{t}Z+\delta}$
gives a well-defined transitive action of $G_{\mathbb{R}}^{0}$ on $\mathcal{H_{S}}$, where $G_{\mathbb{R}}^{0}$ denotes the identity component of $G_{\mathbb{R}}$. The denominator of the above expression is the factor of automorphy
$j(\gamma,Z):=-\frac{1}{2}S_{0}[Z]\gamma+d^{t}Z+\delta.$
Let now $L_{1}:=\mathbb{Z}^{n+4}$ and define the groups
$\Gamma(L_{0}):=\{g\in G_{\mathbb{Q}}^{*}\mid gL_{0}=L_{0}\},$
$\Gamma(L_{1}):=\{g\in G_{\mathbb{Q}}\mid gL_{1}=L_{1}\}.$
Let also $\Gamma^{+}(L_{0})=\Gamma(L_{0})\cap G^{*,0}_{\mathbb{R}}$. If then
$\Gamma_{S}:=G_{\mathbb{R}}^{0}\cap\Gamma(L_{1}),$
we have the following definition:
Definition 2.2.
Let $k\in\mathbb{Z}$. Then a holomorphic function $F:\mathcal{H}_{S}\longrightarrow\mathbb{C}$ is called a modular form of weight $k$ with respect to the group $\Gamma_{S}$ if it satisfies the equation
$(F|_{k}\gamma)(Z):=j(\gamma,Z)^{-k}F(\gamma\langle Z\rangle)=F(Z)$
for all $\gamma\in\Gamma_{S}$ and $Z\in\mathcal{H}_{S}$. We will denote the set of such forms by $M_{k}(\Gamma_{S})$.
Let us now give a couple more definitions on lattices.
Definition 2.3.
Given a $\mathbb{Z}-$lattice $\Lambda$, equipped with a bilinear form $\sigma:V\times V\longrightarrow\mathbb{Q}$, where $V=\Lambda\otimes_{\mathbb{Z}}\mathbb{Q}$, we define its dual lattice by
$\Lambda^{*}:=\{x\in\Lambda\otimes_{\mathbb{Z}}\mathbb{Q}\mid\sigma(x,y)\in\mathbb{Z}\textup{ }\forall y\in\Lambda\}.$
Definition 2.4.
With the notation as above, the level of the lattice $\Lambda$ is given as the least positive integer $q$ such that $\frac{1}{2}q\sigma(x,x)\in\mathbb{Z}$ for every $x\in\Lambda^{*}$.
Now, $F\in M_{k}(\Gamma_{S})$ admits a Fourier expansion of the form (cf. [eisenstein_thesis, page 25]):
$F(Z)=\sum_{r\in L_{0}^{*}}A(r)e(r^{t}S_{0}Z).$
where $Z\in\mathcal{H}_{S}$. It is then Koecher’s principle that gives us that $A(r)=0$ unless $r\in L_{0}^{*}\cap\overline{\mathcal{P_{S}}}$ (cf. [eisenstein_thesis, Theorem 1.5.2]). By [eisenstein_thesis, Theorem 1.6.23], we have the following definition for cusp forms.
Definition 2.5.
$F\in M_{k}(\Gamma_{S})$ is called a cusp form if it admits a Fourier expansion of the form
$F(Z)=\sum_{r\in L_{0}^{*}\cap\mathcal{P_{S}}}A(r)e(r^{t}S_{0}Z).$
We denote the space of cusp forms by $S_{k}(\Gamma_{S})$.
Finally, we give the definition of the Maass space, as this is defined in [krieg_jacobi, Section 5]:
Definition 2.6.
The Maass space $M_{k}^{*}(\Gamma_{S})$ consists of all $F\in M_{k}(\Gamma_{S})$, so that if their Fourier expansion is
$F(Z)=\sum_{r\in L_{0}^{*}}A(r)e(r^{t}S_{0}Z),$
we have
$A(r)=\sum_{d\mid\gcd(\rho)}d^{k-1}A\left(\begin{pmatrix}lm/d^{2}\\ -S^{-1}\lambda/d\\ 1\end{pmatrix}\right),$
where $r=S_{0}^{-1}\rho$, with $\rho=\begin{pmatrix}m\\ \lambda\\ l\end{pmatrix}$.
[…]
For now, assume that $V$ is a vector space of dimension $n<\infty$ over $\mathbb{Q}$ together with a positive definite symmetric bilinear form $\sigma$ and an even lattice $\Lambda$ in $V$, i.e. $\sigma(\lambda,\lambda)\in 2\mathbb{Z}$ for all $\lambda\in\Lambda$. We start with the following definitions:
Definition 3.1.
We define the Heisenberg group to be:
$H^{(\Lambda,\sigma)}(\mathbb{R})=\{(x,y,\zeta)\mid x,y\in\Lambda\otimes\mathbb{R},\zeta\in S^{1}\},$
where $S^{1}:=\{z\in\mathbb{C}\mid|z|=1\}$, equipped with the following composition law:
$(x_{1},y_{1},\zeta_{1})(x_{2},y_{2},\zeta_{2}):=(x_{1}+x_{2},y_{1}+y_{2},\zeta_{1}\zeta_{2}e(\sigma(x_{1},y_{2}))).$
The integral Heisenberg group is defined to be $H^{(\Lambda,\sigma)}(\mathbb{Z}):=\{(x,y,1)\mid x,y\in\Lambda\}$ and in the following we drop the last coordinate for convenience.
Proposition 3.1.
The group $SL_{2}(\mathbb{R})$ acts on $H^{(\Lambda,\sigma)}(\mathbb{R})$ from the right, via:
$((x,y,\zeta),A)\longmapsto(x,y,\zeta)^{A}:=\left((x,y)A,\zeta e\left(\sigma[(x,y)A]-\frac{1}{2}\sigma(x,y)\right)\right).$
where $(x,y)A$ denotes the formal multiplication of the vector $(x,y)$ with $A$, i.e. if $A=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$, we have $(x,y)A:=(ax+cy,bx+dy)$.
Definition 3.2.
The real Jacobi group associated with $(\Lambda,\sigma)$, denoted with $J^{(\Lambda,\sigma)}(\mathbb{R})$, is defined to be the semi-direct product of $SL_{2}(\mathbb{R})$ and $H^{(\Lambda,\sigma)}(\mathbb{R})$. The composition law is then:
$(A,h)\cdot(A^{\prime},h^{\prime}):=(AA^{\prime},h^{A^{\prime}}h^{\prime}).$
We also define the integral Jacobi group to be the semi-direct product of $SL_{2}(\mathbb{Z})$ and $H^{(\Lambda,\sigma)}(\mathbb{Z})$ and we will denote it by $J^{(\Lambda,\sigma)}$.
We are now going to define some slash operators, acting on holomorphic, complex-valued functions on ×$\mathbb{H}\times(\Lambda\otimes\mathbb{C})$.
Definition 3.3.
Let $k$ be a positive integer and $f:\mathbb{H}\times(\Lambda\otimes\mathbb{C})\longrightarrow\mathbb{C}$ a holomorphic function. For $M=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in SL_{2}(\mathbb{R})$ , we define:
$f|_{k,(\Lambda,\sigma)}[M](\tau,z):=(c\tau+d)^{-k}e^{-\pi ic\sigma(z,z)/(c\tau+d)}f\left(\frac{a\tau+b}{c\tau+d},\frac{z}{c\tau+d}\right).$
In the case when $M\in GL_{2}^{+}(\mathbb{R})$, we use $\det(M)^{-1/2}M$ instead of $M$. For $h=(x,y,\zeta)\in H^{(\Lambda,\sigma)}(\mathbb{R})$:
$f\mid_{k,(\Lambda,\sigma)}h(\tau,z):=\zeta\cdot e^{\pi i\tau\sigma(x,x)+2\pi i\sigma(x,z)}f(\tau,z+x\tau+y).$
Finally, for the action of $J^{(\Lambda,\sigma)}(\mathbb{R})$ on complex-valued, holomorphic functions on ×$\mathbb{H}\times(L\otimes\mathbb{C})$, we have:
$(f,(A,h))\longmapsto f\mid_{k,(\Lambda,\sigma)}(A,h)(\tau,z):=(f\mid_{k,(\Lambda,\sigma)}A)\mid_{k,(\Lambda,\sigma)}h.$
We now have the following definition ([jacobi_lattice, Definition 1.23]):
Definition 3.4.
Let $V_{\mathbb{C}}:=V\otimes\mathbb{C}$ and extend $\sigma$ to $V_{\mathbb{C}}$ by $\mathbb{C}$-linearity. For $k$ a positive integer, a holomorphic function $f:\mathbb{H}\times V_{\mathbb{C}}\rightarrow\mathbb{C}$ (where $\mathbb{H}$ denotes the usual upper half plane) is called a Jacobi form of weight $k$ with respect to $(\Lambda,\sigma)$ if the following hold:
  • For all $\gamma\in J^{(\Lambda,\sigma)}$, we have
    $f\mid_{k,(\Lambda,\sigma)}\gamma(\tau,z)=f(\tau,z).$
  • $f$ has a Fourier expansion of the form
    $f(\tau,z)=\sum_{n\in\mathbb{Z},r\in\Lambda^{*},2n\geq\sigma[r]}c_{f}(n,r)e(n\tau+\sigma(r,z)).$
[…]
We now have a notion of a scalar product for elements of $S_{k}(\Lambda,\sigma)$ ([jacobi_lattice, Definition 1.33]).
Definition 3.5.
Let $\phi,\psi\in S_{k}(\Lambda,\sigma)$. If then $U\leq J^{(\Lambda,\sigma)}$ of finite index, we define their Petersson inner product via:
[…]
where $\tau=u+iv$, $z=x+iy$ and $dV:=v^{-n-2}dudvdxdy$. This inner product does not depend on the choice of $U$, so in what follows, we drop the subscript.
From now on, we specify to our case, by taking $\Lambda=L=\mathbb{Z}^{n}$ and $\sigma(x,y)=x^{t}Sy$ for all $x,y\in V$.
[…]
$F(\tau^{\prime}+m,z,\tau)=F(\tau^{\prime},z,\tau)$
(by the form of elements belonging to $\Gamma_{S}$). Hence, we can write
$F(Z)=\sum_{m\geq 0}\phi_{m}(\tau,z)e^{2\pi im\tau^{\prime}},$
[…]
$\phi_{m}(\tau,z)=\sum_{l=0}^{\infty}\sum_{\begin{subarray}{c}\lambda\in\mathbb{Z}^{n},\\ 2ml\geq\lambda^{t}S^{-1}\lambda\end{subarray}}c_{\phi_{m}}(\lambda,l)e^{2\pi i(l\tau+\lambda^{t}z)}.$
We note that then $\phi_{m}\in J_{k}(\mathbb{Z}^{n},m\sigma)$.
We will now re-write the above Fourier expansion of a cusp form in a form similar to Kohnen-Skoruppa in [kohnen_skoruppa]. By the above, we have:
$\phi_{m}(\tau,z)=\sum_{n\in\mathbb{Z},r\in\mathbb{Z}^{n},2mn>r^{t}S^{-1}r}c_{\phi}(n,r)e(n\tau+r^{t}z).$
Now, from the condition $2mn>r^{t}S^{-1}r$ and the definition of the level, we can write
$mnq-\frac{1}{2}qS^{-1}[r]=-D\implies n=\frac{\frac{1}{2}qS^{-1}[r]-D}{qm},$
with some integer $D<0$.
So, in total, we get the following lemma:
Lemma 3.1.
[…]
$\phi_{m}(\tau,z)=\sum_{\begin{subarray}{c}D\in\mathbb{Z}_{<0},r\in\mathbb{Z}^{n}\\ \frac{1}{2}qS^{-1}[r]\equiv D\pmod{qm}\end{subarray}}c_{\phi_{m}}\left(D,r\right)e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{qm}\tau+r^{t}z\right).$
Remark 1.
By adjusting [jacobi_lattice, Proposition 1.25] to the notation of the above Lemma 3.1, we have that
$c_{\phi_{m}}(D,r)=c_{\phi_{m}}(D^{\prime},r^{\prime})\iff D=D^{\prime}\textup{and }S^{-1}r\equiv S^{-1}r^{\prime}\pmod{m\mathbb{Z}^{n}}.$
[…]
Theorem 3.1.
With the notation as above, the mapping $F\longmapsto\phi_{1}$ gives an isomorphism between the Maass space $M_{k}^{*}(\Gamma_{S})$ and $J_{k}(L,\sigma)$, as vector spaces.
4. Poincaré Series
[…]
Definition 4.1.
We define the support of the lattice $L$ with respect to the bilinear form $\sigma(x,y)=x^{t}Sy$ for $x,y\in V$ to be
$\textup{supp}(L,\sigma):=\left\{(D,r)\mid D\in\mathbb{Q}_{\leq 0},r\in L^{*},D\equiv\frac{1}{2}S[r]\pmod{\mathbb{Z}}\right\}.$
Now, if we write $r\longmapsto S^{-1}r$ with $r\in L$, we get
$qD\equiv\frac{1}{2}qS^{-1}[r]\pmod{q\mathbb{Z}},$
which then implies $D\in\frac{1}{q}\mathbb{Z}$. Hence, by writing $D\longmapsto\frac{D}{q}$, we get that equivalently
$\textup{supp}(L,\sigma)=\left\{(D/q,S^{-1}r)\mid D\in\mathbb{Z}_{\leq 0},r\in L,D\equiv\frac{1}{2}qS^{-1}[r]\pmod{q}\right\}.$
Let then
$\widetilde{\textup{supp}}(L,\sigma):=\{(D,r)\mid(D/q,S^{-1}r)\in\textup{supp}(L,\sigma)\}.$
and in the following this is the set we will use.
Definition 4.2.
Let $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$. We then define the following complex valued function on ×$\mathbb{H}\times(L\otimes\mathbb{C})$:
$g_{D,r}(\tau,z):=e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{q}+r^{t}z\right).$
Definition 4.3.
Let $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$ and set
$J^{(L,\sigma)}_{\infty}:=\left\{\left(\begin{pmatrix}1&n\\ 0&1\end{pmatrix},(0,\mu)\right)\mid n\in\mathbb{Z},\mu\in L\right\}.$
The Poincaré series of weight $k$ for the lattice $(L,\sigma)$ is defined by
$P_{k,D,r}(\tau,z):=\sum_{\gamma\in J_{\infty}^{(L,\sigma)}\backslash J^{(L,\sigma)}}g_{D,r}\mid_{k,(L,\sigma)}\gamma(\tau,z).$
We now have the following very important property, which can be found in [jacobi_lattice, Theorem 2.3, (i)].
Proposition 4.1.
Let $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$. Then for any $f\in J_{k}(L,\sigma)$ with
$f(\tau,z)=\sum_{\begin{subarray}{c}D^{\prime}\in\mathbb{Z}_{<0},r^{\prime}\in L\\ \frac{1}{2}qS^{-1}[r^{\prime}]\equiv D^{\prime}\pmod{q}\end{subarray}}c_{f}\left(D^{\prime},r^{\prime}\right)e\left(\frac{\frac{1}{2}qS^{-1}[r^{\prime}]-D^{\prime}}{q}\tau+(r^{\prime})^{t}z\right),$
we have
$\langle f,P_{k,D,r}\rangle=\lambda_{k,D}c_{f}(D,r),$
for some constant $\lambda_{k,D}\in\mathbb{C}$ depending on $k$ and $D$.
5. Operator $V_{N}$ and Dirichlet Series
In this section, we again denote by $\sigma(x,y)=x^{t}Sy$ for $x,y\in V$. Given a Jacobi form $\phi\in J_{k}(\mathbb{Z}^{n},\sigma)$ and $N\geq 1$, we define the operator:
$V_{N}:J_{k}(\mathbb{Z}^{n},\sigma)\longrightarrow J_{k}(\mathbb{Z}^{n},N\sigma),$
given by (compare also with [sugano_lifting, Section 6], [jacobi_lattice, Definition 4.25]):
$(V_{N}\phi)(\tau,z)=N^{k-1}\sum_{M\in SL_{2}(\mathbb{Z})\backslash M_{2}(\mathbb{Z})_{N}}(c\tau+d)^{-k}e^{-\pi icNS[z]/(c\tau+d)}\phi\left(M\langle\tau\rangle,\frac{Nz}{c\tau+d}\right),$
where $M_{2}(\mathbb{Z})_{N}=\{g\in M_{2}(\mathbb{Z})|\det g=N\}$ and for $M=\begin{pmatrix}a&b\\ c&d\end{pmatrix},\tau\in\mathbb{H}$, we define
$M\langle\tau\rangle:=\frac{a\tau+b}{c\tau+d}.$
[…]
Proposition 5.1.
[…]
$\sum_{\begin{subarray}{c}D\in\mathbb{Z}_{<0},r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r]\mod qN\end{subarray}}c_{\phi}\left(D,r\right)e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{qN}\tau+r^{t}z\right)\longmapsto$
$\longmapsto\sum_{\begin{subarray}{c}D<0,r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r]\mod q\end{subarray}}\left(\sum_{d\mid N}d^{k-(n+1)}\sum_{\begin{subarray}{c}s\mod dS\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[s]\mod{qd}\end{subarray}}c_{\phi}\left(\left(\frac{N}{d}\right)^{2}D,\frac{N}{d}s\right)\right)e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{q}\tau+r^{t}z\right).$
Remark 2.
The $(D,r)$ coefficient of $V_{N}^{*}\phi$ is independent of $r$.
Proof.
We start by writing
$V_{N}\phi=N^{k/2-1}\sum_{A\in\Gamma_{1}\backslash M_{2}(\mathbb{Z})_{N}}\phi_{\sqrt{N}}|_{k,(\mathbb{Z}^{n},N\sigma)}\left(\frac{1}{\sqrt{N}}A\right),$
where we define $\phi_{\sqrt{N}}(\tau,z):=\phi\left(\tau,\sqrt{N}z\right)$. We can then follow the proof in [kohnen_skoruppa, pages 554-556] line by line and arrive at the result that the adjoint $V_{N}^{*}:J_{k}(\mathbb{Z}^{n},N\sigma)\longrightarrow J_{k}(\mathbb{Z}^{n},\sigma)$ is given by:
$\phi\longmapsto N^{k/2-2n-1}\sum_{X\mod{N\mathbb{Z}^{2n}}}\sum_{A\in\Gamma_{1}\backslash M_{2}(\mathbb{Z})_{N}}\phi_{\sqrt{N}^{-1}}|_{k,(\mathbb{Z}^{n},N\sigma)}\left(\frac{1}{\sqrt{N}}A\right)|_{k,(\mathbb{Z}^{n},N\sigma)}X.$
Let us now compute the action on the Fourier coefficients. We choose a set of representatives of the form $\begin{pmatrix}a&b\\ 0&d\end{pmatrix}$ with $ad=N$ and $0\leq b<d$. We then obtain that the above expression can be written as:
$N^{k/2-2n-1}\sum_{\lambda,\mu\in\mathbb{Z}^{n}/N\mathbb{Z}^{n}}\sum_{\begin{subarray}{c}ad=N\\ b\in\mathbb{Z}/d\mathbb{Z}\end{subarray}}\left(\frac{d}{\sqrt{N}}\right)^{-k}\phi\left(\frac{a\tau+b}{d},\frac{z+\lambda\tau+\mu}{d}\right)e^{\pi i\tau S[\lambda]+2\pi i\lambda^{t}Sz}.$
By then substituting the Fourier expansion:
$\phi(\tau,z)=\sum_{\begin{subarray}{c}D\in\mathbb{Z}_{<0},r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r]\mod qN\end{subarray}}c_{\phi}\left(D,r\right)e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{qN}\tau+r^{t}z\right),$
for $\phi$, the above can be written as:
$N^{k/2-2n-1}\sum_{\lambda,\mu\in\mathbb{Z}^{n}/N\mathbb{Z}^{n}}\sum_{\begin{subarray}{c}ad=N\\ b\in\mathbb{Z}/d\mathbb{Z}\end{subarray}}\left(\frac{d}{\sqrt{N}}\right)^{-k}\sum_{D,r}c_{\phi}(D,r)e\left(\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{qN}\cdot\frac{a}{d}+\frac{S[\lambda]}{2}+\frac{r^{t}\lambda}{d}\right)\tau+\right.\\ \left.+\left(\frac{r^{t}z}{d}+\lambda^{t}Sz\right)+\frac{\frac{1}{2}qS^{-1}[r]-D}{qN}\cdot\frac{b}{d}+\frac{r^{t}\mu}{d}\right).$
Now, the term
$\sum_{\begin{subarray}{c}b\in\mathbb{Z}/d\mathbb{Z}\\ \mu\in\mathbb{Z}^{n}/N\mathbb{Z}^{n}\end{subarray}}e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{qN}\cdot\frac{b}{d}+\frac{r^{t}\mu}{d}\right)$
is zero unless $d\mid\frac{\frac{1}{2}qS^{-1}[r]-D}{qN}$ and $d\mid r$, meaning it divides all of its components. In that case, the sum equals $dN^{n}$. The conditions then imply that we can replace $r\longmapsto dr$ and $D\longmapsto Dd^{2}$. The last one follows from the fact that $N=ad$ and so $d\mid N$ as well. Hence, we obtain that the expression equals
$N^{k-n-1}\sum_{\lambda\in\mathbb{Z}^{n}/N\mathbb{Z}^{n}}\sum_{d\mid N}d^{1-k}\sum_{\begin{subarray}{c}D<0,r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r]\mod{\frac{qN}{d}}\end{subarray}}c_{\phi}(d^{2}D,dr)e\left(\left(\frac{\frac{1}{2}qS^{-1}[r]+\frac{1}{2}qS[\lambda]+qr^{t}\lambda-D}{q}\right)\tau+\right.\\ \left.+r^{t}z+\lambda^{t}Sz\right)=$
$=N^{k-n-1}\sum_{d\mid N}d^{1-k}\sum_{\lambda\in\mathbb{Z}^{n}/N\mathbb{Z}^{n}}\sum_{\begin{subarray}{c}D<0,r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r-S\lambda]\mod\frac{qN}{d}\end{subarray}}c_{\phi}(d^{2}D,d(r-S\lambda))e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{q}\tau+r^{t}z\right),$
after setting $r\longmapsto r-S\lambda$. Now, as in [kohnen_skoruppa], we set
$\lambda\equiv t+\frac{N}{d}t^{\prime}\mod N,$
with $t\mod\frac{N}{d}$ and $t^{\prime}\mod d$. We then have
$S^{-1}(d(r-S\lambda))\equiv S^{-1}(d(r-St))\mod N\mathbb{Z}^{n}$
and
$D\equiv\frac{1}{2}qS^{-1}[r-St]\mod\frac{qN}{d}.$
The first property can be seen easily and the second follows from the fact that we already have $D\equiv\frac{1}{2}qS^{-1}[r-S\lambda]\mod\frac{qN}{d}$ and also we can check:
$\frac{1}{2}qS^{-1}[r-St]-\frac{1}{2}qS^{-1}[r-S\lambda]=\frac{1}{2}q[2r^{t}(\lambda-t)+S[t]-S[\lambda]].$
So, it suffices to show that
$\frac{1}{2}S[t]\equiv\frac{1}{2}S[\lambda]\mod{\frac{N}{d}}.$
But we can write $\lambda=t+\frac{N}{d}\alpha$ with $\alpha\in\mathbb{Z}^{n}$ and we can then see the claim follows.
Hence, because of Remark 1, the final expression we obtain has the form (after replacing $d$ by $N/d$):
$\sum_{d\mid N}d^{k-(n+1)}\sum_{t\in\mathbb{Z}^{n}/d\mathbb{Z}^{n}}\sum_{\begin{subarray}{c}D<0,r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r-St]\mod qd\end{subarray}}c_{\phi}\left(\left(\frac{N}{d}\right)^{2}D,\frac{N}{d}(r-St)\right)e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{q}\tau+r^{t}z\right).$
What remains to prove now (after fixing $D,r$ with the appropriate conditions) is
$\sum_{d\mid N}d^{k-(n+1)}\sum_{\begin{subarray}{c}t\in\mathbb{Z}^{n}/d\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r-St]\mod qd\end{subarray}}c_{\phi}\left(\left(\frac{N}{d}\right)^{2}D,\frac{N}{d}(r-St)\right)=$
$=\sum_{d\mid N}d^{k-(n+1)}\sum_{\begin{subarray}{c}t\mod dS\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[t]\mod{qd}\end{subarray}}c_{\phi}\left(\left(\frac{N}{d}\right)^{2}D,\frac{N}{d}t\right).$
This follows by setting $u=r-St$ and then observing that $D\equiv\frac{1}{2}qS^{-1}[u]\mod qd$,
$r-St\equiv r-St^{\prime}\mod dS\mathbb{Z}^{n}\iff t\equiv t^{\prime}\mod d\mathbb{Z}^{n}$
and by using the fact that $c_{\phi}(D,s)=c_{\phi}(D,s^{\prime})$ if $s\equiv s^{\prime}\mod NS\mathbb{Z}^{n}$. So we can consider the entries $r-St$ $\mod NS\mathbb{Z}^{n}$ and all of these are different $\mod dS\mathbb{Z}^{n}$. ∎
[…]
$\mathcal{P}_{k,D,r}(\tau^{\prime},z,\tau):=\sum_{N\geq 1}V_{N}P_{k,D,r}(\tau,z)e(N\tau^{\prime}),$
which is then a cusp form of weight $k$ with respect to $\Gamma_{S}$ (see [jacobi_lattice, Theorem 2.3 (i)]) and an element of the Maass space (cf. [sugano_lifting, Lemma 6.1] and Remark 3). Here, by abusing notation, we use $P_{k,D,r}$ to actually denote the Poincaré series $\frac{1}{\lambda_{k,D}}P_{k,D,r}$, with the quantities defined in Section 4. We then define the Dirichlet series:
$\mathcal{D}_{F,\mathcal{P}_{k,D,r}}(s):=\sum_{N=1}^{\infty}\langle\phi_{N},V_{N}P_{k,D,r}\rangle N^{-s}=\sum_{N=1}^{\infty}\langle V_{N}^{*}\phi_{N},P_{k,D,r}\rangle N^{-s}.$
Remark 3.
A corollary of [jacobi_lattice, Theorem 2.3 (i)] is that $S_{k}(L,\sigma)$ is generated by $\{P_{k,D,r}\mid(D,r)\in\widetilde{\textup{supp}}(L,\sigma)\}$ (cf. [jacobi_lattice, Corollary 2.4]). This observation, together with Theorem 3.1, gives us that the Maass space $M_{k}^{*}(\Gamma_{S})$, is generated by $\{\mathcal{P}_{k,D,r}\mid(D,r)\in\widetilde{\textup{supp}}(L,\sigma)\}$.
Let us now relate the coefficients $c_{\phi_{N}}\left(\frac{N^{2}}{d^{2}}D,\frac{N}{d}s\right)$ with the Fourier coefficients of $F$. $F$ admits a Fourier expansion of the form
$F(Z)=\sum_{\tilde{r}\in L_{0}^{*}}A(\tilde{r})e(r^{t}S_{0}Z)=\sum_{N=1}^{\infty}\phi_{N}(\tau,z)e(N\tau^{\prime}),$
where $Z=(\tau^{\prime},z,\tau)$ and $\tilde{r}=(n,r,N)$ with $r\in L^{*}$. We can then write
$\phi_{N}(\tau,z)=\sum_{n\in\mathbb{Z},r\in L^{*}}A(n,r,N)e(n\tau-r^{t}Sz)=\sum_{n\in\mathbb{Z},r\in\mathbb{Z}^{n}}A(n,S^{-1}r,N)e(n\tau-r^{t}z).$
But now $\begin{pmatrix}1_{2}&&\\ &-1_{n}&\\ &&1_{2}\end{pmatrix}\in\Gamma_{S}$ and therefore if $Z=(\tau^{\prime},z,\tau)\in\mathcal{H}_{S}$, we have
$F\left((\tau^{\prime},-z,\tau)\right)=F((\tau^{\prime},z,\tau)),$
which then implies $A(n,r,N)=A(n,-r,N)$ for all $n,N\in\mathbb{Z},r\in L^{*}$. Therefore, after setting $r\longmapsto-r$, we can re-write the above as:
$\phi_{N}(\tau,z)=\sum_{n\in\mathbb{Z},r\in\mathbb{Z}^{n}}A(n,S^{-1}r,N)e(n\tau+r^{t}z).$
Comparing now the above expression with the expression
$\sum_{D,r}c_{\phi_{N}}\left(D,r\right)e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{qN}\tau+r^{t}z\right),$
we see that we after setting $r\longmapsto\frac{N}{d}s,D\longmapsto\frac{N^{2}}{d^{2}}D$, we obtain
$c_{\phi_{N}}\left(\frac{N^{2}}{d^{2}}D,\frac{N}{d}s\right)=A\left(\frac{N}{d}\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)\right).$
From the above considerations, we arrive at the following proposition:
Proposition 5.2.
With the notation as above and $N\geq 1$ we have:
$\langle V_{N}^{*}\phi_{N},P_{k,D,r}\rangle=\sum_{d\mid N}d^{k-(n+1)}\sum_{\begin{subarray}{c}s\mod dS\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[s]\mod{qd}\end{subarray}}A\left(\frac{N}{d}\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)\right).$
Proof.
The proof follows from Proposition 5.1 and Proposition 4.1. ∎
6. Relation to the class number
In this section, our goal is to bring $\mathcal{D}_{F,\mathcal{P}_{k,D,r}}$ into a form similar to the one in [kohnen_skoruppa, page 553]. We first need some definitions of the adelized groups and of genus and classes of lattices.
Let $V$ denote any finite dimensional vector space over $\mathbb{Q}$ of dimension $n\geq 1$. For each prime $p$ (including infinity), we define $V_{p}:=V\otimes_{\mathbb{Q}}\mathbb{Q}_{p}$. For a $\mathbb{Z}-$lattice $L$ in $V$, we denote by $L_{p}:=L\otimes_{\mathbb{Z}}\mathbb{Z}_{p}$. This coincides with the $\mathbb{Z}_{p}-$linear span of $L$ in $V_{p}$.
Proposition 6.1.
Let $L$ be a fixed $\mathbb{Z}-$lattice in $V$. Then, the following are true:
  • If $M$ is another $\mathbb{Z}-$lattice, then $L_{p}=M_{p}$ for almost all $p$. Moreover, $L\subset M$ if $L_{p}\subset M_{p}$ for all $p$ and $L=M$ if $L_{p}=M_{p}$ for all $p$.
  • For all $p<\infty$, let $N_{p}\in V_{p}$ denote a $\mathbb{Z}_{p}-$lattice such that $N_{p}=L_{p}$ for almost all $p<\infty$. Then, there is a $\mathbb{Z}-$lattice $M$ in $V$ such that $M_{p}=N_{p}$ for all $p<\infty$.
Proof.
See [shimura2004arithmetic, Lemma 9.2]. ∎
Definition 6.1.
We define the adelizations $V$ and $\textup{GL}(V)$, as follows:
$V_{\mathbb{A}}=\left\{v\in\prod_{p\leq\infty}V_{p}\mid v_{p}\in L_{p}\textup{ for almost all }p<\infty\right\}$
$\textup{GL}(V)_{\mathbb{A}}=\left\{\alpha\in\prod_{p\leq\infty}\textup{GL}(V_{p})\mid\alpha_{p}L_{p}=L_{p}\textup{ for almost all }p<\infty\right\}$
Here $L$ is an arbitrary $\mathbb{Z}-$lattice, and the above definitions do not depend on the choice of it, by virtue of Proposition 6.1.
Remark 4.
When we identify $V$ with $\mathbb{Q}^{n}$, we identify $\textup{GL}(V)$ with $\textup{GL}_{n}(\mathbb{Q})$. Then, we identify $\textup{GL}_{n}(\mathbb{Q})_{\mathbb{A}}$ with $\textup{GL}_{n}(\mathbb{Q}_{\mathbb{A}})$, where $\mathbb{Q}_{\mathbb{A}}$ is the usual ring of adeles.
Remark 5.
Given $x\in\textup{GL}(V)_{\mathbb{A}}$, we have that $x_{p}L_{p}=L_{p}$ for almost all $p<\infty$. By Proposition 6.1, there is a $\mathbb{Z}-$lattice $M$ such that $M_{p}=x_{p}L_{p}$ for all $p$. We denote this lattice by $xL$. Hence, $xL$ is the lattice which has the property $(xL)_{p}=x_{p}L_{p}$ for all $p<\infty$.
Definition 6.2.
Let $G\subset GL_{n}(\mathbb{Q})$ any algebraic linear group. For any field $K$ containing $\mathbb{Q}$, we denote by $G_{K}$ the group of $K-$rational points in $G$. Now, for any prime number $p$, we abbreviate by $G_{p}$ the group $G_{\mathbb{Q}_{p}}$. Let also $G_{\infty}$ denote $G_{\mathbb{R}}$. We then define
$G_{\mathbb{A}}:=\left\{x\in\prod_{p\leq\infty}G_{p}\mid x_{p}L_{p}=L_{p}\textup{ for almost all }p\right\}.$
Moreover, we use the notation $G_{\mathbb{A},f}$ to denote the finite part of $G_{\mathbb{A}}$.
Definition 6.3.
Let $L$ denote a $\mathbb{Z}-$lattice in $V$ and $G\subset GL_{n}(\mathbb{Q})$ any algebraic group. Then, for any $x\in G$, $xL$ is also a $\mathbb{Z}-$lattice in $V$. We define the class of $L$ to be $\{xL\mid x\in G\}$. Similarly, for any $x\in G_{\mathbb{A}}$, $xL$ is again a $\mathbb{Z}-$lattice in $V$. The set $\{xL\mid x\in G_{\mathbb{A}}\}$ is called the genus of $L$.
Now, the genus of $L$ can be decomposed into disjoint union of classes. If $C=\{x\in G_{\mathbb{A}}\mid xL=L\}$, the map $xC\longmapsto xL$ gives a bijection between $G_{\mathbb{A}}/C$ and the genus of $L$, so gives a bijection between $G\backslash G_{\mathbb{A}}/C$ and the set of classes contained in the genus of $L$. In general, if $U$ is an open subgroup of $G_{\mathbb{A}}$, we call the number #$\#(G\backslash G_{\mathbb{A}}/U)$ the class number of $G$ with respect to $U$.
Now, with all the main definitions out of the way, we want to deal with the vectors appearing in Proposition 5.2. We start with the next lemma:
Lemma 6.1.
Let $D\in\mathbb{Z}_{\leq 0}$ and define the set
$\Xi_{d}:=\left\{\xi=\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}\mid s\mod dS\mathbb{Z}^{n},D\equiv\frac{1}{2}qS^{-1}[s]\mod{qd}\right\}.$
We then have
$\Xi_{d}\subset L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$
for all $d\geq 1$ coprime to $D$. Here, for a $\mathbb{Z}-$lattice $\Lambda$ in $V_{0}$, $\mathfrak{b}$ a fractional ideal of $\mathbb{Q}$ and $q\in\mathbb{Q}^{\times}$, we define:
$\Lambda[q,\mathfrak{b}]=\{x\in V_{0}\mid\phi_{0}[x]=q\textup{ and }\phi_{0}(x,\Lambda)=\mathfrak{b}\}.$
Proof.
Firstly, for any vector $\xi\in\Xi_{d}$, we can directly compute $\phi_{0}[\xi]=-\frac{D}{q}$. It now remains to show that the $\mathbb{Z}-$ideal
$\{\xi^{t}S_{0}x\mid x\in\mathbb{Z}^{n+2}\}$
equals $\mathbb{Z}$ (because the bilinear form is $\phi_{0}=\frac{1}{2}S_{0}$). But for any basis vector $e_{i}$ of the lattice $\mathbb{Z}^{n+2}$, we have
$\xi^{t}S_{0}e_{i}=\begin{cases}d&\textup{if }i=1\\ -s_{i-1}&\textup{if }2\leq i\leq n+1\\ \frac{\frac{1}{2}qS^{-1}[s]-D}{qd}&\textup{if }i=n+2\end{cases}$
so the above ideal is contained in $\mathbb{Z}$ by the conditions that define the set $\Xi_{d}$. Now, if it were equal to $k\mathbb{Z}$ for some $k\geq 1$, then we would have $k\mid d$, $k\mid s$ and $k\mid\frac{\frac{1}{2}qS^{-1}[s]-D}{qd}$ which would then imply $k\mid D$, and so $k=1$ as $d,D$ are assumed to be coprime. ∎
We now have the following very important lemma:
Lemma 6.2.
The class number of $G_{\mathbb{Q}}^{*}$, defined as #$\#(G^{*}_{\mathbb{Q}}\backslash G^{*}_{\mathbb{A}}/C)$, where
$C=\{x\in G^{*}_{\mathbb{A}}\mid xL_{0}=L_{0}\},$
is 1.
Proof.
The proof follows from [shimura2004arithmetic, Theorem 9.26], as the ideal class group of $\mathbb{Q}$ is trivial and $G^{*}_{\mathbb{R}}$ is not compact because we are in the indefinite case (the signature is $(1,n+1)$). Hence, $Y$ in that theorem is the whole $\mathbb{Q}_{\mathbb{A}}^{\times}$ and so $J_{0}$, hence $J_{\phi}$, is the whole $J$, in the notation of the theorem. Also $n\geq 3$ there corresponds to $n\geq 1$ in our notation. ∎
Fix now an element $\xi\in L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$ and consider the algebraic subgroup of $G_{\mathbb{Q}}^{*}$:
$H(\xi)_{\mathbb{Q}}=\left\{g\in G_{\mathbb{Q}}^{*}\mid g\xi=\xi\right\}.$
We note that $H(\xi)_{\mathbb{Q}}=SO^{\psi}(W)$, where $W:=\{x\in V_{0}\mid\phi_{0}(x,\xi)=0\}$ and $\psi:=\phi_{0}\mid_{W}$.
We now have the following Proposition, which is a special case of [shimura_diophantine_2006, Theorem 2.2]:
Proposition 6.2.
There is a bijection
$L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})\longleftrightarrow H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C),$
which is given as follows:
If $k\in L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$, then there is some $\alpha\in G^{*}_{\mathbb{Q}}$ such that $k=\alpha\xi$ (Witt’s theorem, [shimura2004arithmetic, Lemma 1.5 (ii)]). We then assign the $H(\xi)_{\mathbb{Q}}-$class of $\alpha^{-1}L_{0}$ to $k$. In particular, this then gives
#$\#\left(L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})\right)=\#(H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)).$
Proof.
First of all, $\Gamma(L_{0})$ acts on $L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$. This can be seen because if $\gamma\in\Gamma(L_{0}),x\in L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$, then $\phi_{0}[\gamma x]=\phi_{0}[x]=-D/q$ and $\phi_{0}(\gamma x,L_{0})=\phi_{0}(\gamma x,\gamma L_{0})=\phi_{0}(x,L_{0})=\frac{1}{2}\mathbb{Z}$.
The remaining assertions can be seen from [shimura_diophantine_2006, (2.7)]. In the notation there, $D$ is the set $C$ of Lemma 6.2 and therefore we can take $Y$ to be the trivial set in [shimura_diophantine_2006, Theorem 2.2, (iii)]. This then gives exactly the result claimed, because $G_{\mathbb{Q}}^{*}\cap D=G_{\mathbb{Q}}^{*}\cap C=\Gamma(L_{0})$ of Section 2. ∎
Let us now write
$L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]=\bigsqcup_{i=1}^{h}\Gamma(L_{0})\xi_{i}.$
with some $\xi_{i}\in L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$. This in particular implies that $\xi_{i}\in L_{0}^{*}$ for all $i$.
We now assert we can take $\xi_{i}\in\mathcal{P}_{S}=\{y=(y_{1},y,y_{2})\in\mathbb{R}^{n+2}\mid y_{1}>0,\phi_{0}[y]>0\}$ of Section 2 for all $i$. The second condition is clear, as $S_{0}[\xi_{i}]=-2D/q>0$ because we take $D<0$. For the first one, we can always mutliply with $-1_{n+2}\in\Gamma(L_{0})$ and the assertion follows.
Let now $F\in S_{k}(\Gamma_{S})$. For an element $\gamma^{\prime}\in\Gamma_{S}$ of the form $\gamma^{\prime}=\begin{pmatrix}1&0&0\\ 0&\gamma&0\\ 0&0&1\end{pmatrix}$ we have $F(\gamma^{\prime}\langle Z\rangle)=F(\gamma Z)$.
In particular this then implies that $A(\gamma r)=A(r)$ for all $r\in L_{0}^{*}$ and $\gamma\in\Gamma^{+}(L_{0})$, where we use the Fourier expansion of $F$ of Definition 2.5:
$F(Z)=\sum_{r\in L_{0}^{*}\cap\mathcal{P_{S}}}A(r)e(r^{t}S_{0}Z).$
Let us now consider the subgroup of elements $\gamma\in\Gamma(L_{0})$ such that $\begin{pmatrix}1&0&0\\ 0&\gamma&0\\ 0&0&1\end{pmatrix}\in\Gamma_{S}$. The only extra condition we need is that $\gamma\in G^{*,0}_{\mathbb{R}}$ (the connected component of the identity) also.
By [krieg_integral_orthogonal, page 180], we see that if $\gamma=\begin{pmatrix}*&*&*\\ x&*&y\end{pmatrix}$, the condition we need is $x+y>0$. Hence, for any $\gamma\in\Gamma(L_{0})$,
$\begin{pmatrix}1&0&0\\ 0&\gamma&0\\ 0&0&1\end{pmatrix}\in\Gamma_{S}\textup{ or }\begin{pmatrix}1&0&0\\ 0&-\gamma&0\\ 0&0&1\end{pmatrix}\in\Gamma_{S}.$
In the second case, we obtain for $\xi\in\Xi_{d}$ of Lemma 6.1, that:
$A\left(\xi\right)=A(\gamma\xi_{i})=A\left(-\xi_{i}\right)=0,$
because $A(r)=0$ unless $r\in L_{0}^{*}\cap\mathcal{P}_{S}$ (cf. Definition 2.5).
Define now the following quantities:
$n(\xi_{i};d):=\#\left\{s\in\mathbb{Z}^{n}/dS\mathbb{Z}^{n}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd}\textup{ and }\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\gamma\in\Gamma^{+}(L_{0})\right\},$
$m(\xi_{i};d):=\#\left\{s\in\mathbb{Z}^{n}/dS\mathbb{Z}^{n}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd}\textup{ and }\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\Gamma(L_{0})\backslash\Gamma^{+}(L_{0})\right\}.$
From the above considerations, Lemma 6.1 and Proposition 5.2, we may write, for $(N,D)=1$:
$\langle V_{N}^{*}\phi_{N},P_{k,D,r}\rangle=\sum_{i=1}^{h}\sum_{0<d|N}d^{k-(n+1)}\left[n(\xi_{i};d)A\left(\frac{N}{d}\xi_{i}\right)+m(\xi_{i};d)A\left(-\frac{N}{d}\xi_{i}\right)\right]=$
$=\sum_{i=1}^{h}\sum_{0<d|N}d^{k-(n+1)}n(\xi_{i};d)A\left(\frac{N}{d}\xi_{i}\right).$
In particular, we arrive at the following proposition:
Proposition 6.3.
Let $\mathcal{P}$ be a sufficiently large finite set of primes, which includes the prime factors of $D$. Let also
$\zeta_{\xi_{i},\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}n(\xi_{i};N)N^{-s}.$
Define also:
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}\langle V_{N}^{*}\phi_{N},P_{k,D,r}\rangle N^{-s},$
for $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$. We then have that:
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=\sum_{i=1}^{h}\zeta_{\xi_{i}}(s-k+n+1)D_{F,\xi_{i},\mathcal{P}}(s),$
where $D_{F,\xi_{i},\mathcal{P}}(s)=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}A(N\xi_{i})N^{-s}$.
7. Relation to Sugano’s Theorem
In order to now obtain an Euler product, we will make use of the main theorem of Sugano in his paper [sugano]. We first need some setup:
For each prime number $p<\infty$, we define $K_{p}:=G_{p}\cap\textup{SL}_{n+4}(\mathbb{Z}_{p})=G(\mathbb{Z}_{p})$ and let $K_{f}:=\prod_{p<\infty}K_{p}$.
We remind ourselves here that $G_{\infty}^{0}$ acts transitively on $\mathcal{H}_{S}$ (cf. Section 2). Let $\mathcal{Z}_{0}$ denote any point of $\mathcal{H}_{S}$ with real part 00 0 and denote by $K_{\infty}$ its stabiliser in $G_{\infty}^{0}$. Then, we have that $G_{\infty}^{0}/K_{\infty}\cong\mathcal{H}_{S}$.
Definition 7.1.
Let $k\geq 0$. A function $\bm{F}:G_{\mathbb{A}}\longrightarrow\mathbb{C}$ is called a holomorphic cusp form of weight $k$ with respect to $K_{f}$ if the following conditions hold:
  1. (1)
    $\bm{F}(\gamma gu)=\bm{F}(g)\textup{ }\forall\gamma\in G_{\mathbb{Q}},u\in K_{f}$.
  2. (2)
    For any $g=g_{\infty}g_{f}$ with $g_{\infty}\in G_{\infty}^{0}$ and $g_{f}\in G_{\mathbb{A},f}$, $\bm{F}(g_{\infty}g_{f})j(g_{\infty},\mathcal{Z}_{0})^{k}$ depends only on $g_{f}$ and $\mathcal{Z}=g_{\infty}\langle\mathcal{Z}_{0}\rangle$ and is holomorphic on $\mathcal{H}_{S}$ as a function of $\mathcal{Z}$.
  3. (3)
    $\bm{F}$ is bounded on $G_{\mathbb{A}}$.
Denote the above space by $\mathfrak{S}_{k}(K_{f})$.
For each $g_{f}\in G_{\mathbb{A},f}$ and $\mathcal{Z}\in\mathcal{H}_{S}$, we define
$\bm{F}(g_{f};\mathcal{Z})=\bm{F}(g_{\infty}g_{f})j(g_{\infty},\mathcal{Z}_{0})^{k},$
where $g_{\infty}\in G^{0}_{\infty}$ is chosen so that $\mathcal{Z}=g_{\infty}\langle\mathcal{Z}_{0}\rangle$. Let now
$\Gamma(g_{f})=G_{\mathbb{Q}}\cap(G_{\infty}^{0}\times g_{f}K_{f}g_{f}^{-1}),$
which is a discrete subgroup of $G_{\infty}^{0}$. We then have
$\bm{F}(g_{f};\gamma\langle\mathcal{Z}\rangle)=j(\gamma,\mathcal{Z})^{k}\bm{F}(g_{f};\mathcal{Z}).$
for all $\gamma\in\Gamma(g_{f})$ and $\mathcal{Z}\in\mathcal{H}_{S}$. Now, if $X\in V_{0}$, define the element $\gamma_{X}\in G$ by:
$\gamma_{X}:=\begin{pmatrix}1&-X^{t}S_{0}&-\frac{1}{2}S_{0}[X]\\ 0&1_{q}&X\\ 0&0&1\end{pmatrix}.$
Now, the holomorphic function $\bm{F}(g_{f};\mathcal{Z})$ is invariant under $\mathcal{Z}\longmapsto\mathcal{Z}+X$ for $X$ in the lattice:
$L(g_{f}):=\{X\in V_{0}\mid\gamma_{X}\in\Gamma(g_{f})\}.$
Hence, every such function then admits a Fourier expansion of the form:
$\bm{F}(g_{f};\mathcal{Z})=\sum_{\begin{subarray}{c}r\in\hat{L}(g_{f})\\ ir\in\mathcal{H}_{S}\end{subarray}}a(g_{f};r)e[2\phi_{0}(r,Z)],$
where $\hat{L}(g_{f}):=\left\{X\in V_{0}\mid 2\phi_{0}(X,Y)\in\mathbb{Z}\textup{ for all }Y\in L(g_{f})\right\}$ is the dual lattice of $L(g_{f})$.
Finally, let us introduce adelic Fourier coefficients. Let $\chi=\prod_{p\leq\infty}\chi_{p}$ be a character of $\mathbb{Q}_{\mathbb{A}}$ such that $\chi_{\mid\mathbb{Q}}=1$ and $\chi_{\infty}(x)=e[x]$ for all $x\in\mathbb{R}$. For $\eta\in V_{0}$ and $g\in G_{\mathbb{A}}$, we define:
[…]
Now, for $g_{\infty}\in G_{\infty}^{0}$ and $g_{f}\in G_{\mathbb{A},f}$ we have
$\bm{F}_{\chi}(g_{\infty}g_{f};\eta)=a(g_{f};\eta)j(g_{\infty};\mathcal{Z}_{0})^{-k}e[2\phi_{0}(\eta,g_{\infty}\langle\mathcal{Z}_{0}\rangle)].$
We then have from the above definitions (see [sugano, (1.14)]):
  • $\bm{F}_{\chi}(\gamma_{X}gu;\eta)=\chi(2\phi_{0}(\eta,X))\bm{F}_{\chi}(g;\eta)\textup{ for all }X\in V_{\mathbb{A}},u\in K_{f}$,
  • $\bm{F}_{\chi}\left(\begin{pmatrix}\alpha&&\\ &\beta&\\ &&\alpha^{-1}\end{pmatrix}g;\eta\right)=\bm{F}_{\chi}(g;\beta^{-1}\eta\alpha),\textup{ for all }\alpha\in\mathbb{Q}^{\times},\beta\in G^{*}_{\mathbb{Q}}$.
  • $\bm{F}(\gamma_{X}g)=\sum_{\eta\in V_{0}}\bm{F}_{\chi}(g;\eta)\chi(2\phi_{0}(\eta,X))\textup{ for all }X\in V_{0,\mathbb{A}}$.
From [shimura2004arithmetic, Lemma 17.6, (i)], there is a bijection between $\mathfrak{S}(K_{f})$ and $S_{k}(\Gamma_{S})$, which is given by $\bm{F}\longmapsto F(\mathcal{Z}):=\bm{F}(g_{\infty})j(g_{\infty},\mathcal{Z}_{0})^{k}$, where $g_{\infty}\in G_{\infty}^{0}$ is chosen so that $g_{\infty}\langle\mathcal{Z}_{0}\rangle=\mathcal{Z}$. This follows by the fact that the class number of $G_{\mathbb{Q}}$ is 11 11, which we can show in the same way as in Lemma 6.2.
In particular, it then follows that we can specify our $F\in S_{k}(\Gamma_{S})$ of previous sections with $\bm{F}(id,\mathcal{Z})$. In that case, i.e. when $g_{f}=(id,id,\cdots,id)$, which we will just denote by $id$, we have that $L(g_{f})=L=\mathbb{Z}^{n+2}$. Now, $a(id,r)=A(r)$ for all $r\in L_{0}^{*}$, so in particular:
$a(id,\xi_{i})=A(\xi_{i}),$
for all $i$, where $\xi_{i}$’s are the representatives for $L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})$.
Let now $\{u_{i}\}_{i=1}^{h}$ denote a complete system of representatives for $H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$ such that $u_{i,\infty}=1$ for all $i=1,\cdots,h$. We then prove the following lemma:
Lemma 7.1.
Let $\xi$ be a specific element of $L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$, as we have specified it right after Lemma 6.2. If now $\xi_{i}$’s and $u_{i}$’s are as above, we have
$a(u_{i},\xi)=a(id,\xi_{i}),$
for all $i=1,\cdots,h$. Therefore, $a(u_{i},\xi)=A(\xi_{i})$ for all $i=1,\cdots,h$.
Proof.
By the definition of the $\xi_{i}$’s and Witt’s theorem ([shimura2004arithmetic, Lemma 1.5 (ii)]), there is $\alpha\in G_{\mathbb{Q}}^{*}$ such that $\xi_{i}=\alpha\xi$. By the correspondence given by Shimura in Proposition 6.2, we get $\alpha^{-1}L_{0}=u_{i}L_{0}$. This then implies $\alpha u_{i}L_{0}=L_{0}$, so $\alpha u_{i}\in C$ (we remind ourselves that $C=\{x\in G^{*}_{\mathbb{A}}\mid xL_{0}=L_{0}\}$). By definition, we then get $\alpha u_{i}\in K_{f}$. Hence, we obtain:
$\bm{F}_{\chi}(u_{i};\xi)=\bm{F}_{\chi}(u_{i};\alpha^{-1}\xi_{i})=\bm{F}_{\chi}\left(\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix}u_{i};\xi_{i}\right)=\bm{F}_{\chi}\left(\left(\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix},id\right);\xi_{i}\right),$
where $\left(\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix},id\right)$ denotes the element of $G_{\mathbb{A}}$ with infinity part $\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix}$. The second equality above follows from the second bullet point and the third equality from the first bullet point above. By the property of $\bm{F}_{\chi}$, just before the bullet points, we obtain:
$a(u_{i};\xi)e[2\phi_{0}(\xi,Z_{0})]=a(id,\xi_{i})j\left(\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix},Z_{0}\right)^{-k}e\left[2\phi_{0}\left(\xi_{i},\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix}\langle Z_{0}\rangle\right)\right].$
But $j\left(\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix},Z_{0}\right)=1$ and $\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix}\langle Z_{0}\rangle=\alpha Z_{0}$ which then gives that the right hand side above equals:
$a(id;\xi)e[2\phi_{0}(\xi_{i},\alpha Z_{0})]=a(id;\xi)e[2\phi_{0}(\alpha\xi,\alpha Z_{0})]=a(id;\xi)e[2\phi_{0}(\xi,Z_{0})],$
because $\xi_{i}=\alpha\xi$. This then gives the result. ∎
For each prime $p$, let $M(g_{f};\xi)_{p}=H(\xi)_{p}\cap g_{f}K_{f}g_{f}^{-1}$ and $M(g_{f};\xi)_{f}=\prod_{p}M(g_{f};\xi)_{p}$. Let us now recall the main theorem in the paper of Sugano ([sugano, Theorem 1]):
Theorem 7.1.
Let $g_{f}=(id,\cdots,id)\in G_{\mathbb{A},f}$ be fixed and $\xi\in\hat{L}(g_{f})$ with $i\xi\in\mathcal{H}_{S}$. We identify $F\in S_{k}(\Gamma_{S})$ with $\bm{F}(g_{f};Z)\in\mathfrak{S}_{k}(K_{f})$ as above. Denote by $V(g_{f};\xi)$ the space of functions on $H(\xi)_{\mathbb{A}}$, which are left $H(\xi)_{\mathbb{Q}}$ and right $H(\xi)_{\mathbb{A}}\cap C$ invariant. Let, as above, $\{u_{i}\}_{i=1}^{h}$ denote a complete system of representatives for $H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$ such that $u_{i,\infty}=1$ for all $i=1,\cdots,h$. Let $\mathcal{P}$ be a sufficiently large set of primes such that $\bm{F}$ is a simultaneous eigenfunction of the Hecke pair $\mathcal{H}_{p}=(G_{p},K_{p})$ for all $p\notin\mathcal{P}$ and $f\in V(g_{f};\xi)$ be a simultaneous eigenfunction of the Hecke algebra $\mathcal{H}_{p}^{\prime}:=(H(\xi)_{p},M(g_{f};\xi)_{p})$ for all $p\notin\mathcal{P}$. Then, we have:
$\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}\mu(\xi)^{-1}\sum_{i=1}^{h}A(N\xi_{i})\frac{\overline{f}(u_{i})}{e(\xi)_{i}}N^{-(s+k-\frac{n+2}{2})}=\left(\mu(\xi)^{-1}\sum_{i=1}^{h}A(\xi_{i})\frac{\overline{f}(u_{i})}{e(\xi)_{i}}\right)L_{\mathcal{P}}(F;s)\times$
×$\times L_{\mathcal{P}}\left(\overline{f};s+\frac{1}{2}\right)^{-1}\times(\textup{zeta})^{-1},$
where $(\textup{zeta})=\begin{cases}1&\textup{if }n\textup{ odd}\\ \zeta_{\mathcal{P}}(2s)&\textup{if }n\textup{ even}\end{cases}$.
The definitions of the Hecke algebras can be found in [sugano, Section 2]. $L(-,s)$ denotes the standard $L-$function attached to orthogonal modular forms, as this is defined in [sugano, Paragraph 4-1]. Also, $\zeta(s)$ denotes the Riemann zeta function and for any $L-$function, we write $L_{\mathcal{P}}$ for the Euler product not containing the primes in $\mathcal{P}$. Also, we define $e(\xi)_{i}=\#\{H(\xi)_{\mathbb{Q}}\cap M(u_{i}g_{f};\xi)_{f}\}$ for $1\leq i\leq h$ and $\mu(\xi)=\sum_{i=1}^{h}e(\xi)^{-1}_{i}$.
Remark 6.
We have adapted the above theorem in our setting, meaning we have taken $g_{f}=(id,id,\cdots,id)$ and then substituted $a(u_{i};N\xi)$ with $A(N\xi_{i})$ in the original form of [sugano, Theorem 1]. We also note that $H(\xi)_{\infty}M(g_{f};\xi)_{f}=H(\xi)_{\mathbb{A}}\cap C$.
By setting then $g(u_{i})=\frac{\overline{f}(u_{i})}{e(\xi)_{i}}\mu(\xi)^{-1}$ and $A_{g}=\sum_{i=1}^{h}g(u_{i})A(\xi_{i})$, the formula in Theorem 7.1 becomes:
$(\textup{zeta})\times L_{\mathcal{P}}\left(\overline{f};s+\frac{1}{2}\right)\sum_{i=1}^{h}g(u_{i})D_{F,\xi_{i},\mathcal{P}}\left(s+k-\frac{n+2}{2}\right)=A_{g}L_{\mathcal{P}}(F;s),$
where $D_{F,\xi_{i},\mathcal{P}}$ is the Dirichlet series appearing in Proposition 6.3. This is true for any simultaneous eigenfunction $f$ (hence $g$ as well) of the Hecke algebras $H_{p}^{\prime}=(H(\xi)_{p},M(g_{f},\xi)_{f})$ for all $p\notin\mathcal{P}$. Our aim is to try to invert it so that we solve for $D_{F,\xi_{i},\mathcal{P}}$.
Now, any such $f$ is a $\mathbb{C}-$valued function on $H(\xi)_{\mathbb{A}}$, which is left $H(\xi)_{\mathbb{Q}}$ invariant and right $H(\xi)_{\infty}M(g_{f};\xi)_{f}$ invariant. By what we are given, we have that
$H(\xi)_{\mathbb{A}}=\bigsqcup_{i=1}^{h}H(\xi)_{\mathbb{Q}}u_{i}D,$
where $D=H(\xi)_{\infty}M(g_{f};\xi)_{f}$. We note $D$ is an open subgroup of $H(\xi)_{\mathbb{A}}$ and $D\cap H(\xi)_{f}$ is compact. By [shimura2004arithmetic, Lemma 17.6], there is a correspondence
$f\longleftrightarrow\{f^{(i)}|i=1,\cdots,h\},$
with each $f^{(i)}\in\mathbb{C}$, because $H(\xi)_{\infty}\cong SO(n+1)$ is compact. We also note here that $f(u_{i})=f^{(i)}$ for all $i=1,\cdots,h$, as we can see by the way these $f^{(i)}$ are defined in the proof of [Shimura_Euler_Product, Lemma 10.8].
Now, for any two simultaneous eigenfunctions $f_{i},f_{j}$ of the Hecke algebras defined by the pairs $\mathcal{H}_{p}^{\prime}=(H(\xi)_{p},M(g_{f},\xi)_{p})$ for all $p\notin\mathcal{P}$, their inner product is defined via the formula:
$\langle f_{i},f_{j}\rangle=\left\{\sum_{k=1}^{h}\nu\left(\Gamma^{k}\right)\right\}^{-1}\sum_{k=1}^{h}\nu\left(\Gamma^{k}\right)\overline{f_{i}^{(k)}}f_{j}^{(k)}=\left\{\sum_{k=1}^{h}\nu\left(\Gamma^{k}\right)\right\}^{-1}\sum_{k=1}^{h}\nu\left(\Gamma^{k}\right)\overline{f_{i}(u_{k})}f_{j}(u_{k}),$
where $\Gamma^{k}=H(\xi)_{\mathbb{Q}}\cap u_{i}Du_{i}^{-1}$ and $\nu\left(\Gamma^{k}\right)=\#(\Gamma^{k})^{-1}$, as in [shimura2004arithmetic, (17.23)] (here we again use the fact that $H(\xi)_{\infty}$ is compact).
As $e(\xi)_{i}=\#\{H(\xi)_{\mathbb{Q}}\cap M(u_{i}g_{f};\xi)\}$, we have that $e(\xi)_{i}=\nu\left(\Gamma^{i}\right)^{-1}$, which also gives
$\mu(\xi)=\sum_{k=1}^{h}\nu\left(\Gamma^{k}\right).$
It is now possible to choose a basis of orthogonal Hecke eigenforms $\{f_{1},f_{2},\cdots,f_{h}\}$ (we necessarily have $h$ eigenforms) with respect to the above inner product. This is true, because the Hecke algebra defined by $\mathcal{H}_{p}^{\prime}$ is commutative and consists of self-adjoint operators for all $p\notin\mathcal{P}$ (cf. [shimura2004arithmetic, Proposition 17.14]). Hence, we get the expression:
$D_{F,\xi_{i},\mathcal{P}}\left(s+k-\frac{n+2}{2}\right)=\mu(\xi)^{-1}(\textup{zeta})^{-1}L_{\mathcal{P}}(F;s)\sum_{j=1}^{h}\nu\left(\Gamma^{i}\right)\frac{\mu(\xi)}{e(\xi)_{i}^{-1}}f_{j}(u_{i})L_{\mathcal{P}}\left(\overline{f_{j}};s+\frac{1}{2}\right)^{-1}A_{f_{j}},$
which after the simplifications becomes:
$D_{F,\xi_{i},\mathcal{P}}(s)=(\textup{zeta})^{-1}\left(s-k+\frac{n+2}{2}\right)L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\sum_{j=1}^{h}f_{j}(u_{i})L_{\mathcal{P}}\left(\overline{f_{j}};s-k+\frac{n+3}{2}\right)^{-1}A_{f_{j}}.$
We note $A_{f_{j}}$ is $A_{g}$ with $g(u_{i})=\frac{\overline{f_{j}(u_{i})}}{e(\xi)_{i}}\mu(\xi)^{-1}$. Hence, we arrive at the following theorem:
Theorem 7.2.
Let $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$. Let $\mathcal{P}$ be a finite set of primes, containing the primes described in [sugano, Theorem 1] and all the prime divisors of $D$. Let $F\in S_{k}(\Gamma_{S})$ be a simultaneous eigenfunction for the Hecke algebra $\mathcal{H}_{p}$, defined by the pair $(G_{p},K_{p})$ for all $p\notin\mathcal{P}$ and let $P_{k,D,r}$ denote the Poincaré series of Section 4. Let also $\{u_{i}\}_{i=1}^{h}$ denote a complete system of representatives for $H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$ such that $u_{i,\infty}=1$ for all $i=1,\cdots,h$ and $\{\xi_{i}\}_{i=1}^{h}$ representatives of $L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$ corresponding to $u_{i}$, in the sense described in Proposition 6.2. Let also $\{f_{j}\}_{j=1}^{h}$ denote a complete orthogonal basis of simultaneous eigenfunctions for the pairs $\mathcal{H}_{p}^{\prime}=(H(\xi)_{p},M(g_{f};\xi)_{p})$ for all $p\notin\mathcal{P}$, and denote with $L_{\mathcal{P}}(-,s)$ the standard $L-$function attached to either $F$ or any $f_{j}$, by ignoring the $p-$factors for $p\in\mathcal{P}$. Finally, let $\zeta_{\xi_{i},\mathcal{P}}(s)$ be as in Proposition 6.3. We then have:
$\mathcal{D}_{F,\mathcal{P}_{K,D,r},\mathcal{P}}(s)=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}\langle V_{N}^{*}\phi_{N},P_{k,D,r}\rangle N^{-s}=$
$=L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\sum_{j=1}^{h}A_{f_{j}}L_{\mathcal{P}}\left(\overline{f_{j}};s-k+\frac{n+3}{2}\right)^{-1}\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)f_{j}(u_{i})\times$
×$\times\begin{cases}1&\textup{if }n\textup{ odd}\\ \zeta_{\mathcal{P}}(2s-2k+n+2)^{-1}&\textup{if }n\textup{ even}\end{cases},$
where $A_{g}=\sum_{i=1}^{h}g(u_{i})A(\xi_{i})$, with $g(u_{i})=\frac{\overline{f_{j}(u_{i})}}{e(\xi)_{i}}\mu(\xi)^{-1}$.
Proof.
By substituting the expression we deduced above in Proposition 6.3, we obtain (here we denote by zeta the function of Theorem 7.1 after $s\longmapsto s-k+(n+2)/2$)
$\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}\langle V_{N}^{*}\phi_{N},P_{k,D,r}\rangle N^{-s}=\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)D_{F,\xi_{i},\mathcal{P}}(s)=$
$=(\textup{zeta})^{-1}\times\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\sum_{j=1}^{h}f_{j}(u_{i})L_{\mathcal{P}}\left(\overline{f_{j}};s-k+\frac{n+3}{2}\right)^{-1}A_{f_{j}}=$
$=(\textup{zeta})^{-1}\times L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\sum_{j=1}^{h}A_{f_{j}}L_{\mathcal{P}}\left(\overline{f_{j}};s-k+\frac{n+3}{2}\right)^{-1}\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)f_{j}(u_{i}).$
Hence, we would like to explore the connection between $\sum_{i=1}^{h}\zeta_{\xi_{i}}(s-k+n+1)f_{j}(u_{i})$ and $L\left(\overline{f_{j}};s-k+\frac{n+3}{2}\right)$.
It turns out we can now obtain a clear-cut Euler product expression in the case $h=1$ and when $D$ is a specifically chosen number. The question is how we can pick $S$, so that we can get $h=1$. This is the theme of the next section.
8. Explicit Examples
We will now focus our attention to some specific examples of matrices $S$ and corresponding Poincaré series. In particular, we will choose matrices $S$, so that the number of representatives for $L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})$ is 11 11 and consider the inner product with the Poincaré series $\mathcal{P}_{k,-q,r}$, with $r\in L$ such that $(-q,r)\in\widetilde{\textup{supp}}(L,\sigma)$. In particular, those choices imply that we can take
$\xi=(1,0,\cdots,0,1)^{t}$
as an element of $L_{0}\left[1,\frac{1}{2}\mathbb{Z}\right]$ (because we choose $D=-q$). The main reason we make these choices is the following remark:
Remark 7.
[…]
$\gamma\xi=\begin{pmatrix}*&*&*\\ x&*&y\end{pmatrix}\xi_{0}=\begin{pmatrix}*\\ x+y\end{pmatrix}$
and then $x+y=d>0$. Hence, by the above, $\gamma\in\Gamma(L_{0})\cap G^{*,0}(\mathbb{R})$. Therefore, with the above choices, $\zeta_{\xi,\mathcal{P}}(s)$ of Proposition 6.3 can be written as
$\zeta_{\xi,\mathcal{P}}(s)=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}n(\xi;N)N^{-s},$
where this time
$n(\xi;d)=\#\left\{s\in\mathbb{Z}^{n}/dS\mathbb{Z}^{n}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd}\right\}.$
In these cases, we are able to deduce an exact Euler product expression, connecting the Dirichlet series of interest and the standard $L-$function of the orthogonal group.
8.1. Examples with rank 11 11
We consider the case where $S=2t$ for some $t\geq 1$ with $t$ square-free. This condition is needed so that the lattice $L=\mathbb{Z}$ (and therefore $L_{0}$ and $L_{1}$) is maximal(cf. [eisenstein_thesis, Example 1.6.61.6.6 1.6.61.6.6(ii)]). Now $V_{0}=\mathbb{Q}^{3}$ and the quadratic form of interest is then:
$\phi_{0}(x,y)=\frac{1}{2}x^{t}S_{0}y,$
for $x,y\in V_{0}$, where $S_{0}=\begin{pmatrix}&&1\\ &-2t&\\ 1&&\end{pmatrix}$. Hence, $\phi_{0}$ is represented by $\frac{1}{2}S_{0}$ with respect to the standard basis $e_{1},e_{2},e_{3}$ of $V$.
By [Shimura_Euler_Product, Paragraph 7.3] we have that there exists a quaternion algebra $B$ over $\mathbb{Q}$ such that we can put $V_{0}=B^{\circ}\zeta$, $\phi_{0}[x\zeta]=dxx^{\iota}$, $2\phi_{0}(x\zeta,y\zeta)=dTr_{B/\mathbb{Q}}(xy^{\iota})$, where $B^{\circ}=\{x\in B\mid x^{\iota}=-x\}$ with $\iota$ the main involution of $B$ and $\zeta\in A(V_{0})$ such that $\zeta^{2}=-d$. Here, by $A(V_{0})$ we mean the Clifford algebra of $(V_{0},\phi_{0})$ (see [shimura2004arithmetic, Section 2]). Also, in general, we define $Tr_{B/\mathbb{Q}}(x):=x+x^{\iota}$ for $x\in B$.
Now, [Shimura_Euler_Product, Paragraph 7.3] gives us a way to compute $\zeta$ and $d$. We first need a basis $h_{1},h_{2},h_{3}$ of $V$ such that $\phi_{0}(h_{i},h_{j})=c_{i}\delta_{ij}$ for all $1\leq i,j\leq 3$. In our case, we can make a choice:
$h_{1}=e_{1}+e_{3},h_{2}=e_{2},h_{3}=e_{1}-e_{3}.$
Then, we get that the condition above is satisfied with $c_{1}=1,c_{2}=-t,c_{3}=-1$ and therefore $d=c_{1}c_{2}c_{3}=t$. Also, even though it is not needed in what follows, $\zeta=h_{1}h_{2}h_{3}$.
We now remind ourselves that $\xi=(1,0,1)^{t}$ and $W=(\mathbb{Q}\xi)^{\perp}$. This then implies that $\phi_{0}[\xi]=1$. From [shimura_diophantine_2006, Paragraph 5.2], we get that $\exists k\in B^{\circ}$ such that $\xi=k\zeta$. Then, if $K:=\mathbb{Q}+\mathbb{Q}k$, we get that $K=\mathbb{Q}(\sqrt{-t})$ because $-t$, which is $-d\phi_{0}[h]$ in the notation there, cannot be a square in $\mathbb{Q}^{\times}$.
We will now use [shimura_diophantine_2006, Theorem 5.7] and the formula $(5.11)$ given there. That, together with the fact that the genus of $L_{0}$ is 11 11 in $V_{0}$ (see Lemma 6.2), will give us the number #$\#\left(L_{0}\left[1,\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})\right)$ by $(5.15\textup{a})$ in [shimura_diophantine_2006, Theorem 5.10]. This will equal $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]$ and we are interested in the cases where this number is 11 11.
We start by first computing the quantities appearing in the statement of [shimura_diophantine_2006, Theorem 5.7]. We have that $\mathfrak{e}$ is the product of all the prime ideals in $\mathbb{Q}$ for which $B$ ramifies. However, for each prime $p$, $\phi_{0}$, viewed as a bilinear form over $\mathbb{Q}_{p}$ is isotropic (the Witt index has to be 11 11 for all primes $p$). Therefore by [Shimura_Euler_Product, Paragraph 7.3], $B$ over $\mathbb{Q}_{p}$ is not a division algebra, hence is isomorphic to $M_{2}(\mathbb{Q}_{p})$, i.e. $B$ splits or is unramified over $p$. Hence, $\mathfrak{e}=\mathbb{Z}$.
Now, $\mathfrak{a}=t\mathbb{Z}$ because $t$ is square-free and for the different $\mathfrak{d}_{K/F}$ we have
$\mathfrak{d}_{K/F}=\begin{cases}2\sqrt{-t}\mathcal{O}_{K}&\textup{ if }-t\nequiv 1\pmod{4}\\ \sqrt{-t}\mathcal{O}_{K}&\textup{ if }-t\equiv 1\pmod{4}.\end{cases}$
We now want to investigate the condition $\mathfrak{c}_{p}^{2}N_{K/F}(\mathfrak{d}_{K/F})_{p}=\mathfrak{a}_{p}\phi_{0}[\xi]\phi_{0}(\xi,L_{0})_{p}^{-2}$ for a prime $p$ for the conductor $\mathfrak{c}$. But, by the above:
$N_{K/F}(\mathfrak{d}_{K/F})=\begin{cases}4t\mathbb{Z}&\textup{ if }-t\nequiv 1\pmod{4}\\ t\mathbb{Z}&\textup{ if }-t\equiv 1\pmod{4},\end{cases}$
and $\phi_{0}[\xi]=1$, $\phi_{0}(\xi,L_{0})=\frac{1}{2}\mathbb{Z}$ and so we get
$\mathfrak{c}_{p}^{2}\cdot\begin{cases}4t\mathbb{Z}_{p}&\textup{ if }-t\nequiv 1\pmod{4}\\ t\mathbb{Z}_{p}&\textup{ if }-t\equiv 1\pmod{4}\end{cases}=4t\mathbb{Z}_{p}.$
Therefore, if $p\neq 2$, $\mathfrak{c}_{p}=\mathbb{Z}_{p}$ and if $p=2$, $\mathfrak{c}_{2}=\mathbb{Z}_{2}$ if $-t\nequiv 1\pmod{4}$ and $\mathfrak{c}_{2}=2\mathbb{Z}_{2}$ otherwise.
Also, $N_{K/F}(\mathcal{O}_{K}^{\times})=\{1\}$ because $K$ is imaginary quadratic and so [×$\left[\mathbb{Z}^{\times}:N_{K/F}(\mathcal{O}_{K}^{\times})\right]=2$. Finally, $\infty$ ramifies and so $\nu=1$. We will now consider specific cases in order to determine the ideal $\mathfrak{a}^{*}$ and the sets $U,U^{\prime}$.
  • $t=1$. Then, $\mathfrak{a}=\mathfrak{e}=\mathfrak{a}^{*}=\mathbb{Z}$ and from the above $\mathfrak{c}=\mathbb{Z}$ as well. Also, $U=\mathcal{O}_{K}^{\times}=\{\pm 1,\pm i\}$ and now
    $U^{\prime}=\{x\in\mathcal{O}_{K}^{\times}\mid x-1\in 2\mathcal{O}_{K,p}\textup{ }\forall p\}.$
    Clearly ±$\pm 1\in U^{\prime}$ but ±$\pm i\notin U^{\prime}$ because $2=(1+i)^{2}=(1-i)^{2}$. So, $[U:U^{\prime}]=2$. Finally, $\mu=0$ because $\mathfrak{a}^{*}\mathfrak{e}=\mathbb{Z}$ and therefore, we get
    $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=(1/1)\cdot 2^{1-0-1}\cdot 2\cdot 2^{-1}=1,$
    as claimed.
  • $t=2$. Then $\mathfrak{a}=2\mathbb{Z}$ and we can compute $\mathfrak{a}^{*}=2\mathbb{Z}$. This is true because the only prime factor of $\mathfrak{a}$ is 22 22 and $r_{2}(\xi)=-2\in\frac{1}{2}\mathbb{Z}_{p}$ and $r_{2}(\xi)\notin\mathfrak{C}_{2}$ (if $-2=u^{2}+4w$ with $u,v\in\mathbb{Z}_{2}$, then $2\mid u$ which then implies $2\mid 1$, contradiciton). The definitions of $r_{2}(\xi)$ and $\mathfrak{C}$ can be found in [shimura_diophantine_2006, Paragraph 5.5]. Also, by the above, $\mathfrak{c}=\mathbb{Z}$. Now, $U=\mathcal{O}_{K}^{\times}=\{\pm 1\}$ and
    $U^{\prime}=\{x\in\mathcal{O}_{K}^{\times}\mid x-1\in 2\sqrt{-2}\mathcal{O}_{K,p}\textup{ }\forall p\neq 2\}.$
    But both ±$\pm 1$ belong to $U^{\prime}$ as $\sqrt{-2}$ is a unit in $\mathcal{O}_{K,p}$ for all $p\neq 2$. Hence, $[U:U^{\prime}]=1$. Finally, $\mu=1$ in this case because 2=($2=(\sqrt{-2})^{2}$ is ramified. Therefore:
    $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=(1/1)\cdot 2^{1-1-1}\cdot 2\cdot 1^{-1}=1.$
  • $t=3$. Here $\mathfrak{a}=3\mathbb{Z}$ and $\mathfrak{a}^{*}=3\mathbb{Z}$ because $r_{3}(\xi)=-\frac{4}{3}\in\frac{1}{3}\mathbb{Z}_{3}$ but $r_{3}(\xi)\notin\mathfrak{C}_{3}$ as $\mathfrak{C}_{3}=\mathbb{Z}_{3}$. By what we said above for the conductor, we get $\mathfrak{c}_{2}=2\mathbb{Z}_{2}$ and $\mathfrak{c}_{p}=\mathbb{Z}_{p}$ for $p\neq 2$. Hence, $\mathfrak{c}=2\mathbb{Z}$. In this case, $\mathcal{O}_{K}^{\times}=\{\pm 1,\pm\omega,\pm\omega^{2}\}$, where $\omega=\frac{1}{2}(1+\sqrt{-3})$. Then,
    $U^{\prime}=\{x\in\mathcal{O}_{K}^{\times}\mid x-1\in 2\sqrt{-3}\mathcal{O}_{K,p}\textup{ }\forall p\neq 3\}.$
    But for $p\neq 3$, $\sqrt{-3}$ is a unit in $\mathcal{O}_{K,p}$. So, the condition becomes $x-1\in 2\mathcal{O}_{K,p}\textup{ }\forall p\neq 3$. We can then check that this is true only for ±$\pm 1\in U$. Therefore, $[U:U^{\prime}]=3$. Also, $\mu=1$ in this case, because 33 33 ramifies in $K$. Also, as $-3\equiv 5\pmod{8}$, we have that 22 22 remains prime in $K$, hence
    $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=(1/1)\cdot 2^{1-1-1}\cdot 2\cdot 3^{-1}\cdot 2\cdot(1+1/2)=1.$
  • Assume $t>3$ with $-t\nequiv 1\pmod{4}$. We write $t=p_{1}\cdots p_{k}$ with $p_{i}$ prime factors. In this case, $\mathfrak{e}=\mathfrak{c}=\mathbb{Z}$ and $\mathfrak{a}=t\mathbb{Z}$. For all $p_{i}$, we have that $r_{p_{i}}(\xi)$ satisfies the conditions of [shimura_diophantine_2006, Theorem 5.7] and therefore $\mathfrak{a}^{*}=t\mathbb{Z}$. We have $\mu=k$ as each $p_{i}$ is ramified in $K$. Now $U=\mathcal{O}_{K}^{\times}=\{\pm 1\}$ and
    $U^{\prime}=\{x\in\mathcal{O}_{K}^{\times}\mid x-1\in 2\sqrt{-t}\mathcal{O}_{K,p}\textup{ }\forall p\neq p_{i},i=1,\cdots,k\}.$
    But for all $p\neq p_{i}$, $\sqrt{-t}$ is a unit in $\mathcal{O}_{K,p}$ and therefore ±$\pm 1\in U^{\prime}$, i.e. $[U:U^{\prime}]=1$. We then obtain:
    $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=c_{K}\cdot 2^{1-k-1}\cdot 2=c_{K}\cdot 2^{1-k}$
    and therefore this is 11 11 if $c_{K}=2^{k-1}$. Hence, the answer in this case is the number fields $K=\mathbb{Q}(\sqrt{-t})$ that satisfy $c_{K}=2^{k-1}$, where $k$ is the number of prime factors of $t$ and $-t\nequiv 1\pmod{4}$. For example, when $c_{K}=2$, $t$ must have 2 prime factors, and examples would be $t=6,10,$ etc.
  • Finally, assume $t>3$ with $-t\equiv 1\pmod{4}$. We write $t=p_{1}\cdots p_{k}$ with $p_{i}$ primes. In these cases, similarly to the case $t=3$, we have $\mathfrak{a}=t\mathbb{Z}$, $\mathfrak{a}^{*}=t\mathbb{Z}$ and $\mathfrak{c}=2\mathbb{Z}$. Also, $\mu=k$ because each prime $p_{i}$ is ramified in $K$. Now, $U=\mathcal{O}_{K}^{\times}=\{\pm{1}\}$ and then
    $U^{\prime}=\{x\in\mathcal{O}_{K}^{\times}\mid x-1\in 2\sqrt{-t}\mathcal{O}_{K,p}\textup{ }\forall p\neq p_{i},i=1,\cdots,k\}.$
    But for $p\neq p_{i}$, $\sqrt{-t}$ is a unit in $\mathcal{O}_{K,p}$ and therefore ±$\pm 1\in U^{\prime}$, which means $[U:U^{\prime}]=1$. Now, if $-t\equiv 5\pmod{8}$, we get that 22 22 is inert in $K$ and if $-t\equiv 1\pmod{8}$, 22 22 splits in $K$. Therefore
    $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=\begin{cases}c_{K}\cdot 2^{1-k-1}\cdot 2\cdot 2\cdot(1+1/2)=3c_{K}\cdot 2^{1-k}&\textup{ if }-t\equiv 5\pmod{8}\\ c_{K}\cdot 2^{1-k-1}\cdot 2\cdot 2\cdot(1-1/2)=c_{K}\cdot 2^{1-k}&\textup{ if }-t\equiv 1\pmod{8}.\end{cases}$
    In the first case the index cannot be 11 11, while in the second case, we must have $c_{K}=2^{k-1}$. Therefore, the answer in this case is $t$ such that $-t\equiv 1\pmod{8}$ so that if $K=\mathbb{Q}(\sqrt{-t})$ we have $c_{K}=2^{k-1}$, where $k$ is the number of distinct prime factors of $t$. For example, the only example for $k=2$ is $t=15$.
Hence, we arrive at the following proposition:
Proposition 8.1.
Let $t$ be one of the following:
  • $t=1,2,3$.
  • $t\nequiv 3\pmod{4}$ and if $t=p_{1}\cdots p_{k}$, $K:=\mathbb{Q}(\sqrt{-t})$, we have $c_{K}=2^{k-1}$.
  • $t\equiv 7\pmod{8}$ and if $t=p_{1}\cdots p_{k}$, $K:=\mathbb{Q}(\sqrt{-t})$, we have $c_{K}=2^{k-1}$.
Then, with the notation as above, we have $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=1$.
Let now $t$ be one of the above. We aim to give an Euler product expression for $\mathcal{D}_{F,\mathcal{P}_{k,-q,r},\mathcal{P}}$. In particular, from Theorem 7.2, we need to give an Euler product expression for
$\zeta_{\xi,\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}n(\xi;N)N^{-s},$
where now $n(\xi;N)=\#\{s\in\mathbb{Z}/2tN\mathbb{Z}\mid s^{2}\equiv-4t\pmod{4tN}\}$, as we have explained in Remark 7.
Now, as $t$ is square-free, we obtain from $s^{2}\equiv-4t\pmod{4tN}$, that $2t\mid s$, so it suffices to look for the number of solutions of the congruence
$ts^{2}\equiv-1\pmod{N},$
with $s\pmod{N}$. This last number of solutions is multiplicative in $N$, so we can write
$\zeta_{\xi,\mathcal{P}}(s)=\prod_{p\notin\mathcal{P}}\left(\sum_{k=0}^{\infty}n(\xi;p^{k})p^{-ks}\right).$
Now, for all $p\notin\mathcal{P}$, we have $(p,t)=1$, so by [congruences, Proposition 14], we get that
$n(\xi;p^{k})=1+\left(\frac{-t}{p}\right),$
for all $k\geq 1$, where $\left(\frac{\cdot}{p}\right)$ denotes the Legendre symbol. Therefore, if we define $\chi_{t}(p):=\left(\frac{-t}{p}\right)$, we deduce that (bear in mind that $\chi_{t}^{2}=1$):
$\zeta_{\xi,\mathcal{P}}(s)=\zeta_{\mathcal{P}}(s)\zeta_{\mathcal{P}}(2s)^{-1}\zeta_{\mathcal{P}}(s,\chi_{t}).$
Therefore, we have the following theorem:
Theorem 8.1.
Let $S=2t$, with $t$ being chosen as in Proposition 8.1. We define $\chi_{t}(p):=\left(\frac{-t}{p}\right)$ and $\psi(p):=\left(\frac{-1}{p}\right)$. Let also $\mathcal{P}$ a sufficiently large finite set of primes, containing the primes described in Theorem 7.2 and the primes so that the conditions of [shimura_orthogonal, Proposition 5.13] are satisfied. We then have:
$\mathcal{D}_{F,\mathcal{P}_{k,-q,r},\mathcal{P}}=A(\xi)L_{\mathcal{P}}\left(F;s-k+\frac{3}{2}\right)\zeta_{\mathcal{P}}(2s-2k+4)^{-1}\frac{\zeta_{\mathcal{P}}(s-k+2,\chi_{t})}{\zeta_{\mathcal{P}}(s-k+2,\psi)},$
where again $\xi=(1,0,1)^{t}$.
Proof.
The proof follows by Theorem 7.2 after choosing $f$ to be the constant 1 on $H(\xi)_{\mathbb{A}}$. We have computed $\zeta_{\xi,\mathcal{P}}$ above and
$L_{\mathcal{P}}(\textbf{1},s-k+2)=\zeta_{\mathcal{P}}(s-k+2,\psi)\zeta_{\mathcal{P}}(s-k+2),$
as this can be computed by [shimura_orthogonal, Proposition 5.15], because $\mathcal{P}$ is chosen so that conditions of [shimura_orthogonal, Proposition 5.13] are satisfied. ∎
Remark 8.
In the case $t=1$, we recover the result of Kohnen and Skoruppa in [kohnen_skoruppa]. For $t>1$, the result can be interpreted as a result of modular forms for a paramodular group (cf. [krieg_integral_orthogonal, Corollary 6], [paramodular_hecke]).
8.2. The rank $n\geq 2$ case
In the rank $n\geq 2$ case, a Theorem like [shimura_diophantine_2006, Theorem 5.7] is not available. For this reason, we seek examples of matrices $S$ so that the following conditions hold:
  1. (1)
    The lattice $L=\mathbb{Z}^{n}$ is $\mathbb{Z}-$maximal.
  2. (2)
    With the notation as in Section 6, $L_{0}\cap W$ is a $\mathbb{Z}-$maximal lattice in $W$ and if $D=\{\alpha\in H(\xi)_{\mathbb{A}}\mid\alpha(L_{0}\cap W)=L_{0}\cap W\}$, we have $D=H(\xi)_{\mathbb{A}}\cap C$.
  3. (3)
    The number of classes in the genus of maximal lattices (this is independent of the choice of the maximal lattice, see [shimura2004arithmetic, Paragraph 9.7]) in $H(\xi)_{\mathbb{Q}}=SO^{\psi}(W)$ is 11 11. Here, $\psi:=\phi_{0}\mid_{W}$.
We will show that for rank $n=2,4,6,8$, there is at least one positive definite even symmetric matrix $S$ of rank $n$, so that the above conditions are satisfied. The reasons why we only provide these examples is explained in Section 9. We start with the following lemma:
Lemma 8.1.
We have $H(\xi)_{\mathbb{Q}}=SO^{\psi}(W)$ with $\psi$ the bilinear form induced by the matrix:
$T=\frac{1}{2}\begin{pmatrix}-2&\\ &-S\end{pmatrix}.$
Proof.
We have $W=\{x\in V_{0}\mid\phi_{0}(x,\xi)=0\}$. Then, it is easy to see that
$W=\textup{span}\{(1,\textbf{0},-1),(0,1,0,\cdots,0,0),(0,0,1,\cdots,0,0),\cdots,(0,0,0,\cdots,1,0)\}.$
By then evaluating $\phi_{0}[x]$ for the above basis vectors $x$ of $W$, the lemma follows. ∎
We claim the following choices for the matrix $S$ satisfy the conditions $(1)-(3)$ above.
  • $n=2$: $S=\begin{pmatrix}2&-1\\ -1&2\end{pmatrix},\begin{pmatrix}2&-1\\ -1&8\end{pmatrix}$ with determinants $3,15$ respectively.
  • $n=4$: $S=\begin{pmatrix}2&-1&-1&-1\\ -1&2&1&0\\ -1&1&2&0\\ -1&0&0&2\end{pmatrix},\begin{pmatrix}2&-1&0&0\\ -1&2&0&0\\ 0&0&2&-1\\ 0&0&-1&2\end{pmatrix}$ with determinants $5,9$ respectively.
  • $n=6$: $S=\begin{pmatrix}2&1&-1&1&-1&1\\ 1&2&0&1&-1&1\\ -1&0&2&-1&1&0\\ 1&1&-1&2&-1&0\\ -1&-1&1&-1&2&0\\ 1&1&0&0&0&2\end{pmatrix}$ with determinant 33 33.
  • $n=8$: $S=\begin{pmatrix}2&-1&1&1&-1&-1&1&-1\\ -1&2&0&-1&0&1&-1&1\\ 1&0&2&1&-1&0&0&0\\ 1&-1&1&2&-1&-1&0&-1\\ -1&0&-1&-1&2&1&-1&1\\ -1&1&0&-1&1&2&-1&1\\ 1&-1&0&0&-1&-1&2&-1\\ -1&1&0&-1&1&1&-1&2\end{pmatrix}$ with determinant 11 11.
Remark 9.
The matrices $S$ of rank 22 22 correspond to the unitary groups of the imaginary quadratic fields $\mathbb{Q}(\sqrt{-3}),\mathbb{Q}(\sqrt{-15})$ respectively (cf. [eisenstein_thesis, Example 1.6.6, (v)]).
Let us first check the conditions $(1)$ and $(3)$. Condition $(1)$ follows by [eisenstein_thesis, Proposition 1.6.12] for the matrices with square-free determinant and [eisenstein_thesis, Lemma 1.6.5] for the matrix of determinant 99 99. Condition $(3)$ follows by [hanke2011enumerating, Section 8], as $W$ corresponds to the definite orthogonal group induced by the matrices:
  • $n=2$: Examples $4,26$ in matrices of 33 33 variables.
  • $n=4$: Examples $3,5$ in matrices of 55 55 variables.
  • $n=6$: Example 33 33 in matrices of 77 77 variables.
  • $n=8$: Example 11 11 in matrices of 99 99 variables.
All these examples correspond to one class in the genus of the standard lattice $\mathbb{Z}^{n+1}$, in the cases when it is maximal. This can also be seen by a simple computation in MAGMA for example. As we mentioned above, this shows that every maximal lattice has one class in its genus. Finally, we need to check condition $(2)$. We will use [shimura2004arithmetic, Proposition 11.12]. We have the following local result:
Proposition 8.2.
Let $S$ denote any of the matrices above. Let $D_{p}=\{\alpha\in H(\xi)_{p}\mid\alpha(L_{0,p}\cap W_{p})=L_{0,p}\cap W_{p}\}$ for each prime $p$. We then have that $L_{0,p}\cap W_{p}$ is a $\mathbb{Z}_{p}-$maximal lattice in $W_{p}$ and also
$D_{p}=H(\xi)_{p}\cap C_{p},$
where $C_{p}=\{x\in G_{p}\mid xL_{0,p}=L_{0,p}\}$.
Proof.
Our proof is based on [shimura2004arithmetic, Proposition 11.12]. We will first establish the following claim:
$L_{0,p}^{*}\neq L_{0,p}\iff p\mid\det(S)$.
Indeed, assume $p\nmid\det(S)$. Then, as $L_{0,p}\subset L_{0,p}^{*}$ always, suffices to show that $L_{0,p}^{*}\subset L_{0,p}$ too. But $L_{0,p}^{*}=S_{0}^{-1}L_{0,p}$ and so if $y\in L_{0,p}^{*}$, we write $y=S_{0}^{-1}x$ with $x\in L_{0,p}$ and so the result follows as $p\nmid\det(S)$ (we have $S_{0}^{-1}=\begin{pmatrix}&&1\\ &-S^{-1}&\\ 1&&&\end{pmatrix}$).
On the other hand, if $p\mid\det(S)$, we can always find $y\in L_{0,p}^{*}\backslash L_{0,p}$.
We remind ourselves that $\phi_{0}[\xi]=1$ and $\phi_{0}(\xi,L_{0,p})=\frac{1}{2}\mathbb{Z}_{p}$ for all primes $p$. Therefore, $\phi_{0}(\xi,L_{0,p})^{2}=\phi_{0}[\xi]\mathbb{Z}_{p}$ for all $p\neq 2$. Moreover, $L_{0,2}^{*}=L_{0,2}$ for every choice of $S$, as $2\nmid\det(S)$ for any $S$. This means [shimura2004arithmetic, Proposition 11.12] is applicable in every case.
Let now $t_{p}$ denote the dimension of the maximal anisotropic subspace of $(\mathbb{Q}_{p},\phi_{0})$ (cf. [shimura2004arithmetic, Paragraph 8.3]). For $p\nmid\det(S)$, we have $L_{0,p}^{*}=L_{0,p}$, $t_{p}\neq 1$ as $n$ is even and $4\phi_{0}[\xi]^{-1}\phi_{0}(\xi,L_{0,p})^{2}=\mathbb{Z}_{p}$. So, by [shimura2004arithmetic, Proposition 11.12, (iii), (2)], $L_{0,p}\cap W_{p}$ is $\mathbb{Z}_{p}-$maximal.
If now $p\mid\det(S)$, we claim $t_{p}>1$ and therefore [shimura2004arithmetic, Proposition 11.12, (iii), (1)] will be applicable. We define $K_{0}:=\mathbb{Q}_{p}(\sqrt{\delta})$, where $\delta:=(-1)^{(n+2)(n+1)/2}\det(\phi_{0})$. By the proof of [shimura_classification_quadratic_forms, Lemma 3.3], we have $t_{p}=2$ if and only if $K_{0}\neq\mathbb{Q}_{p}$. We show this case by case:
  • $S=\begin{pmatrix}2&-1\\ -1&2\end{pmatrix}$. Then, we claim that $t_{3}=2$. Now, $\det(\phi_{0})=-\frac{3}{2^{4}}$ and then $K_{0}=\mathbb{Q}_{3}(\sqrt{\det(\phi_{0})})=\mathbb{Q}_{3}(\sqrt{-3})\neq\mathbb{Q}_{3}$.
  • $S=\begin{pmatrix}2&-1\\ -1&8\end{pmatrix}$. In this case, $\det(\phi_{0})=-\frac{15}{2^{4}}$ and we claim $t_{3}=t_{5}=2$. But again, if $p\in\{3,5\}$, we have $K_{0}=\mathbb{Q}_{p}(\sqrt{\det{\phi_{0}}})=\mathbb{Q}_{p}(\sqrt{-15})\neq\mathbb{Q}_{p}$ by taking valuations (for example, if $\sqrt{-15}\in\mathbb{Q}_{3}$, then $-15=u^{2}$ for some $u\in\mathbb{Q}_{3}$ and so $1\displaystyle 2v_{3}(u)=$ 1, contradiction).
  • $S=\begin{pmatrix}2&-1&-1&-1\\ -1&2&1&0\\ -1&1&2&0\\ -1&0&0&2\end{pmatrix}$. In this case, $\det(\phi_{0})=-\frac{5}{2^{6}}$ and so $K_{0}=\mathbb{Q}_{5}(\sqrt{5})\neq\mathbb{Q}_{5}$, so $t_{5}=2$.
  • $S=\begin{pmatrix}2&-1&0&0\\ -1&2&0&0\\ 0&0&2&-1\\ 0&0&-1&2\end{pmatrix}$. In this case, $K_{0}=\mathbb{Q}_{3}$ and we can compute $t_{3}=4$ in Sage.
  • $S=\begin{pmatrix}2&1&-1&1&-1&1\\ 1&2&0&1&-1&1\\ -1&0&2&-1&1&0\\ 1&1&-1&2&-1&0\\ -1&-1&1&-1&2&0\\ 1&1&0&0&0&2\end{pmatrix}$. In this case, $\det(\phi_{0})=-\frac{3}{2^{8}}$ and so again $K_{0}=\mathbb{Q}_{3}(\sqrt{-3})\neq\mathbb{Q}_{3}$. Therefore, $t_{3}=2$.
  • In the $n=8$ case, we have that $\det(S)=1$, so there are no primes to check ($L_{0,p}^{*}=L_{0,p}$ for all primes $p$).
Finally, the fact that $D_{p}=H(\xi)_{p}\cap C_{p}$ follows from [shimura2004arithmetic, Proposition 11.12, (iv)], as $t_{p}\neq 1$ always, because $n$ is even. ∎
We are now ready to give the global statement.
Proposition 8.3.
With $S$ be any of the matrices above, we have that $L_{0}\cap W$ is $\mathbb{Z}-$maximal in $W$ and if $D=\{\alpha\in H(\xi)_{\mathbb{A}}\mid\alpha(L_{0}\cap W)=L_{0}\cap W\}$, then $D=H(\xi)_{\mathbb{A}}\cap C$.
Proof.
The first claim follows by [eisenstein_thesis, Proposition 1.6.9], as all the localizations are maximal by Proposition 8.2. For the second one, we have
$H(\xi)_{\mathbb{A}}\cap C=\{x\in H(\xi)_{\mathbb{A}}\mid xL_{0}=L_{0}\}.$
But the lattice $xL_{0}$ is the lattice which is defined by $(xL_{0})_{p}=x_{p}L_{0,p}$ for all primes $p$. Now, if $x\in H(\xi)_{\mathbb{A}}$ with $xL_{0}=L_{0}$, we have $x(L_{0}\cap W)=L_{0}\cap W$ (cf. [shimura2004arithmetic, page 104]), so $H(\xi)_{\mathbb{A}}\cap C\subset D$.
On the other hand, if $x\in D$, then $x_{p}\in D_{p}=H(\xi)_{p}\cap C_{p}$ by Proposition 8.2 and therefore $x_{p}L_{0,p}=L_{0,p}$ for all primes $p$. This means that $x\in H(\xi)_{\mathbb{A}}\cap C$, as wanted. ∎
8.3. Euler product expression for the Dirichlet series
The question of this section is to obtain an Euler product expression for the Dirichlet series of interest in each of the above cases and relate it to standard $L-$function attached to $F$. By Theorem 7.2, the first step is to determine $\zeta_{\xi,\mathcal{P}}(s)$. Hence, we need to compute the quantity:
$n(\xi;d)=\#\left\{s\in\mathbb{Z}^{n}/dS\mathbb{Z}^{n}\mid\frac{1}{2}qS^{-1}[s]\equiv-q\pmod{qd}\right\},$
with $\xi=(1,\mathbf{0},1)^{t}$, as we have explained in Remark 7. The steps we follow are:
  1. (1)
    Find unimodular integer matrices $P,Q$ such that $PSQ=\textup{diag}(a_{1},\cdots,a_{n})$, for some positive integers $a_{i}$.
  2. (2)
    We then substitute $t=Ps\implies s=P^{-1}t$. Then, we have
    $s-s^{\prime}\in dS\mathbb{Z}^{n}\iff t-t^{\prime}\in dPS\mathbb{Z}^{n}\iff t-t^{\prime}\in dPSQ\mathbb{Z}^{n},$
    because $Q$ is unimodular. Hence, if $t=(t_{1},\cdots,t_{n})^{t}$, we consider each $t_{i}\pmod{da_{i}}$.
  3. (3)
    We then solve the congruence
    $\frac{1}{2}qS^{-1}[P^{-1}t]\equiv-q\pmod{qd}.$
Let us now deal with the specific examples we have. In the following, let
$\chi_{S}(p):=\left(\frac{(-1)^{n/2}\det{S}}{p}\right).$
  1. (1)
    $S=\begin{pmatrix}2&-1\\ -1&2\end{pmatrix}$. Here, we have $PSQ=\textup{diag}(1,3)$, with $P=\begin{pmatrix}1&1\\ 1&2\end{pmatrix}$. We then have
    $P^{-1}t=\begin{pmatrix}2t_{1}-t_{2}\\ -t_{1}+t_{2}\end{pmatrix}$
    and after substituting, the congruence of interest becomes ($q=3$ here):
    $3t_{1}^{2}-3t_{1}t_{2}+t_{2}^{2}\equiv-3\pmod{3d}$
    with $t_{1}\pmod{d}$ and $t_{2}\pmod{3d}$. Now, by the form of the equation, we get $3\mid t_{2}$, so the congruence becomes:
    $t_{1}^{2}-3t_{1}t_{2}+3t_{2}^{2}\equiv-1\pmod{d}$
    with $t_{1},t_{2}\pmod{d}$. Now, we have $n(\xi;d)=N(T;d)$, where the last quantity is defined as the number of solutions to the congruence $T[t]\equiv-1\pmod{d}$, with
    $T=\begin{pmatrix}1&-3/2\\ -3/2&3\end{pmatrix}.$
    But $N(T;d)$ is multiplicative in $d$. Let now $p\notin\mathcal{P}\cup\{2,3\}$, so that 22 22 has a multiplicative inverse $\pmod{p}$ and $p\nmid\det{T}$. Then, by [diagonalisation], we know that for each $k\geq 1$, there is a matrix $U_{k}$ such that $T[U_{k}]\equiv R\pmod{p^{k}}$ with some diagonal matrix $R$. By then setting $t\longmapsto U_{k}t$, we can still consider $t_{i}\pmod{p^{k}}$ for all $i$ (as the determinant of $U$ is non-zero $\pmod{p^{k}}$). We then have $N(T;p^{k})=N(R;p^{k})$, with $R$ diagonal. We can now count the number of solutions $N(R;p^{k})$ by [congruences, Proposition 4]. In particular, if $R=\textup{diag}(a_{1},a_{2})$, we have
    $N(R;p^{k})=p^{k}\left[1-\frac{1}{p}\left(\frac{-a_{1}a_{2}}{p}\right)\right],$
    because $c_{1}(-1)=1,c_{p}(-1)=-1,c_{p^{l}}(-1)=0$ for $l\geq 2$ in the notation there. But $\det{R}=a_{1}a_{2}$, and $\det{R}\equiv(\det{U_{k}})^{2}\det{T}\pmod{p}$, so
    $\left(\frac{-a_{1}a_{2}}{p}\right)=\left(\frac{-\det{T}}{p}\right)=\left(\frac{-3u^{2}}{p}\right)=\left(\frac{-3}{p}\right)=\chi_{S}(p),$
    where $2u\equiv 1\pmod{p}$. Therefore, we obtain:
    $n(\xi;p^{k})=N(R;p^{k})=p^{k}\left[1-\frac{\chi_{S}(p)}{p}\right].$
  2. (2)
    $S=\begin{pmatrix}2&-1\\ -1&8\end{pmatrix}$. In this case, we have $PSQ=\textup{diag}(1,15)$ with $P=\begin{pmatrix}1&1\\ 1&2\end{pmatrix}$. By following the above steps, the congruence becomes:
    $15t_{1}^{2}-15t_{1}t_{2}+4t_{2}^{2}\equiv-15\pmod{15d}$
    with $t_{1}\pmod{d}$ and $t_{2}\pmod{15d}$. But now $15\mid t_{2}$ and so after $t_{2}\longmapsto 15t_{2}$, we get
    $t_{1}^{2}-15t_{1}t_{2}+60t_{2}^{2}\equiv-1\pmod{d}$
    with $t_{1},t_{2}\pmod{d}$ and with exact same reasoning as before, we get:
    $n(\xi;p^{k})=p^{k}\left[1-\frac{\chi_{S}(p)}{p}\right].$
    for all $p\notin\mathcal{P}\cup\{2,3,5\}$ with $n(\xi;d)$ multiplicative in $d$.
  3. (3)
    $S=\begin{pmatrix}2&-1&-1&-1\\ -1&2&1&0\\ -1&1&2&0\\ -1&0&0&2\end{pmatrix}$. In this case, we have $PSQ=\textup{diag}(1,1,1,5)$ with
    $P=\begin{pmatrix}3&1&1&3\\ 1&1&0&1\\ 1&0&1&1\\ 3&1&1&4\end{pmatrix}.$
    By then substituting $s=P^{-1}t$ the congruence becomes:
    $5t_{1}^{2}+5t_{2}^{2}+5t_{3}^{2}+2t_{4}^{2}-5t_{1}t_{2}-5t_{1}t_{3}-5t_{1}t_{4}+5t_{2}t_{3}\equiv-5\pmod{5d},$
    with $t_{1},t_{2},t_{3}\pmod{d}$ and $t_{4}\pmod{5d}$. But, we see that $5\mid t_{4}$, so after setting $t_{4}\longmapsto 5t_{4}$, the congruence becomes:
    $t_{1}^{2}+t_{2}^{2}+t_{3}^{2}+10t_{4}^{2}-t_{1}t_{2}-t_{1}t_{3}-5t_{1}t_{4}+t_{2}t_{3}\equiv-1\pmod{d},$
    with $t_{i}\pmod{d}$ for all $i$. As before, this then gives
    $n(\xi;p^{k})=p^{3k}\left[1-\frac{\chi_{S}(p)}{p^{2}}\right].$
    for all $p\notin\mathcal{P}\cup\{2,5\}$ with $n(\xi;d)$ multiplicative in $d$.
  4. (4)
    $S=\begin{pmatrix}2&-1&0&0\\ -1&2&0&0\\ 0&0&2&-1\\ 0&0&-1&2\end{pmatrix}$. In this case, using the same method, we obtain:
    $3t_{1}^{2}+12t_{2}^{2}+t_{3}^{2}+t_{4}^{2}-9t_{1}t_{2}+3t_{1}t_{3}-6t_{2}t_{3}+3t_{2}t_{4}\equiv-3\pmod{3d},$
    with $t_{1},t_{2}\pmod{d}$ and $t_{3},t_{4}\pmod{3d}$. This then implies $3\mid t_{3}^{2}+t_{4}^{2}$ and so $3\mid t_{3},t_{4}$. Therefore, as previously, we obtain
    $n(\xi;p^{k})=p^{3k}\left[1-\frac{\chi_{S}(p)}{p^{2}}\right].$
    for all $p\notin\mathcal{P}\cup\{2,3\}$ with $n(\xi;d)$ multiplicative in $d$.
  5. (5)
    In the last two cases, we omit the calculations, but by applying the exact same reasoning as above, we will again get:
    $n(\xi;p^{k})=p^{k(n-1)}\left[1-\frac{\chi_{S}(p)}{p^{n/2}}\right].$
    for all $p$ not in $\mathcal{P}$, $p\neq 2$ and $p\nmid\det{S}$ and with $n(\xi;d)$ multiplicative in $d$. Here $n=6,8$.
In all of the above cases, we therefore obtain (here $\mathcal{P}$ is possibly enlargened but still finite):
$\zeta_{\xi,\mathcal{P}}(s)=\prod_{p\notin\mathcal{P}}\zeta_{\mathcal{P}}(s-(n-1))\zeta_{\mathcal{P}}(s-n/2+1,\chi_{S})^{-1}.$
We therefore arrive at the following Theorem:
Theorem 8.2.
Let $S$ denote any of the matrices of Section 8.2 with even rank $n$ and $\mathcal{P}$ a sufficiently large finite set of primes, containing the primes described in Theorem 7.2 and the primes so that the conditions of [shimura_orthogonal, Proposition 5.13] are satisfied. Then, if we define
$\chi_{S}(p):=\left(\frac{(-1)^{n/2}\det{S}}{p}\right),$
we have
$\mathcal{D}_{F,\mathcal{P}_{k,-q,r},\mathcal{P}}(s)=A(\xi)\zeta_{\mathcal{P}}(2s-2k+n+2)^{-1}\zeta_{\mathcal{P}}\left(s-k+\frac{n+4}{2},\chi_{S}\right)^{-1}L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\times$
×$\times\prod_{i=1}^{n-1}\zeta_{\mathcal{P}}(s-k+(n+2)-i)^{-1},$
where again $\xi=(1,\mathbf{0},1)^{t}$.
Proof.
The proof follows from Theorem 7.2 after we choose $f$ to be the constant 1 on $H(\xi)_{\mathbb{A}}$. From [shimura_orthogonal, Proposition 5.15], we have that (since $\mathcal{P}$ is chosen so that conditions of [shimura_orthogonal, Proposition 5.13] are satisfied):
$L_{\mathcal{P}}\left(\mathbf{1},s-k+\frac{n+3}{2}\right)=\prod_{i=1}^{n}\zeta_{\mathcal{P}}(s-k+(n+2)-i),$
as in general $L_{\textup{Sug}}\left(\textbf{1};s-\frac{n-1}{2}\right)=L_{\textup{Shi}}(s)$. Here, the notation means the $L-$function we encounter in [sugano] and [shimura_orthogonal] respectively, as these are normalised differently. Finally, the computation of $\zeta_{\xi,\mathcal{P}}$ has been performed above. ∎
9. Conclusions
In this paper, our main goal has been to obtain Euler product expansions for the Dirichlet series $\mathcal{D}_{F,\mathcal{P}_{k,D,r}}(s)$, extending in this way, the method of Kohnen and Skoruppa in [kohnen_skoruppa] to the orthogonal setting. However, it seems that our case presents some limitations, which do not allow us to give uniform results for orthogonal groups of arbitrary rank.
To be more specific, the first main difficulty arises from the fact that it does not seem to exist an obvious connection between $\sum_{i=1}^{h}\zeta_{\xi_{i}}(s-k+n+1)f_{j}(u_{i})$ and $L\left(\overline{f_{j}};s-k+(n+3)/2\right)$. This does not happen in Kohnen and Skoruppa, as this connection can be seen by using the correspondence between equivalence classes of binary quadratic forms with fixed discriminant and ideal classes in the class group of a quadratic extension. This has led us restricting our attention to the case where $h=1$.
In order therefore to find explicit examples, we need to pick $\xi$ such that the size of $H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$ is just 11 11. The choice of $\xi=(1,\textbf{0},1)^{t}$, apart from being natural and simple, satisfies another very important purpose. In particular, it gives us that any vector in $\Xi_{d}$ of Lemma 6.1 lies in the same equivalence class of $\xi$ under not only the action of $\Gamma(L_{0})$, but also of $\Gamma^{+}(L_{0})$. This is crucial for the explicit computation of $\zeta_{\xi,\mathcal{P}}(s)$, which is done in Section 8. In the rank 11 11 case, Shimura in [shimura_diophantine_2006] has given an explicit formula for the index $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]$, which we use to find all the cases for which this index is 11 11 (see Section 8.1). In the rank $n\geq 2$, such a formula is not available, so we need an explicit description of $H(\xi)_{\mathbb{Q}}$ as a definite orthogonal group. It is well-known that there is a finite number of definite orthogonal groups with class number 11 11 and their rank is at most 1010 1010 (cf. [watson]). Using, therefore, the enumeration of definite orthogonal groups with class number 11 11 in [hanke2011enumerating] as well as [shimura2004arithmetic, Proposition 11.12], we arrive at the specific examples of Section 8 for rank $n\geq 2$.
[…]